### v1

# A Multi-Perspective Analysis of Memorization in Large Language Models

Abstract
Large Language Models (LLMs), trained on massive corpora with billions of parameters, show unprecedented performance in various fields. Though surprised by their excellent performances, researchers also noticed some special behaviors of those LLMs. One of those behaviors is memorization, in which LLMs can generate the same content used to train them. Though previous research has discussed memorization, the memorization of LLMs still lacks explanation, especially the cause of memorization and the dynamics of generating them. In this research, we comprehensively discussed memorization from various perspectives and extended the discussion scope to not only just the memorized content but also less and unmemorized content. Through various studies, we found that: (1) Through experiments, we revealed the relation of memorization between model size, continuation size, and context size. Further, we showed how unmemorized sentences transition to memorized sentences. (2) Through embedding analysis, we showed the distribution and decoding dynamics across model size in embedding space for sentences with different memorization scores. The n-gram statistics analysis presents d (3) An analysis over n-gram and entropy decoding dynamics discovered a boundary effect when the model starts to generate memorized sentences or unmemorized sentences. (4)We trained a Transformer model to predict the memorization of different models, showing that it is possible to predict memorizations by context.
1 Introduction
Large language models (LLMs), spanning from BERT Devlin et al. (2019) to the recent strongest GPT-4 OpenAI et al. (2024), have not just revolutionized the Natural Language Processing (NLP) but nearly the whole artificial intelligence research field. Though their performances on various tasks are surprising, we still know little about the underlying mechanism of Large Language Models whose structures are based on neural models known as black-box Alain and Bengio (2018). Due to the enormous model size and pre-train data size, unprecedented phenomena in previous machine learning are also identified in the LLM, while in this research we discuss the memorization.
Memorization Hartmann et al. (2023) in the context of LLMs means the LLM sometimes will verbatim output the same sentence recorded in their pre-training corpus. Memorization is a coin with two sides, on the one hand, utilizing such memorization, we can use LLMs as a knowledge base Petroni et al. (2019) that can be served for multiple usage, on the other hand, personal information contained in the pre-train corpora may also be elicited maliciously Yao et al. (2024). Though such a phenomenon has been utilized or studied since the merge of LLMs, detailed analysis of memorization is still under-explored.
In this research, we discuss memorization in detail and analyze it from several perspectives, e.g., the pre-train data level, the encoding level, the generation level, and finally, we try to predict memorization. Following those studies, we found that:
  • For both memorization and unmemorized sentences, even though their numbers increase/decrease with the model size, the trend is not linear, showing the maximum capacity to memorize sentences exists. The number of memorized sentences decreases sub-linearly with the continuation size while increasing above linearly with the context size. We also study how unmemorized sentences transition to memorized sentences and vice versa. We found that memorized sentences are more likely to be memorized when scaling the model size, while low-memorized sentences will be gradually memorized.
  • Through a study of the encoding dynamics, we found that sentences with different memorization scores are clustered in the embedding space. Additionally, with the generation of tokens, the embedding similarly remains basically stable while the Euclidean distance is shrinking. Moreover, the larger the model, the more significant the difference in the Cosine Similarity. This may help to explain why the larger model performs better than the smaller-size model.
  • In both the n-gram and decoding entropy dynamics, we found boundary phenomenon exists when the model generates memorized or unmemorized sentences. In the n-gram, the frequency when generating unmemorized tokens suddenly decreases while increasing for memorized tokens. In the entropy analysis, the entropy suddenly increases when generation unmemorized sentences while decreasing for memorized sentences.
  • Finally, we trained a naive Transformer model to predict memorization based on context. We show that prediction accuracy over tokens is high and grows with the model size, while it is still hard to predict correctly for a whole sentence
2 Related Works
2.1 Scaling Laws of LLM
In this research, the conducted experiments span across various model sizes, thus relating to the research of Scaling Laws Kaplan et al. (2020); Abnar et al. (2021); Villalobos (2023), which suggests the LLM’s performance scales with the size of corpora, the parameter of model and the computation required.
Those researches inspired LLMs to continuously scale those factors to gain higher performance like T5 Raffel et al. (2023), GPT-3 Brown et al. (2020), PaLM 2 Anil et al. (2023), etc. On the other hand, researchers also analyzed how LLM Scaling affects particular tasks rather than general performance. For example, Sun and Miceli-Barone (2024) discussed how LLM scaling affects both translation tasks and prompt injection attacks. Within those discussions, the LLM’s emergent abilities Wei et al. (2022) are also discovered, which means the LLM suddenly reaches high performance on previous low-performance tasks when the LLM reaches a specific model size such as digital addition Nye et al. (2021), reasoning Lampinen et al. (2022) etc. Recent studies have questioned whether emergent abilities are just mirages caused by misuse of metrics Schaeffer et al. (2023) or whether emergent abilities are just context-learning Lu et al. (2023).
In the field of memorization, Carlini et al. (2023) discussed the memorization across model size and found that the number of memorized texts grows with the model size and context size. Additionally, Biderman et al. (2023a) also discussed similar topics and found that a large portion of memorized text in a small-size model is also memorized by a larger model, showing that memorized texts may share certain common features.
2.2 Memorization
Memorization, which means the neural model tends to memorize the case during training, is a long-existing phenomenon in machine learning. While the emergence of LLMs recalled this topic.
Before LLM, a phenomenon close to memorization in real-world practices is known as over-fitting Tetko et al. (1995) which means a near-zero loss in the train set while bad performance in test set. Zhang et al. (2017) described that the neural model could reach near-zero loss on the training set, which suggests the neural model perfectly memorized the relation between input and data label. However, memorization is different from over-fitting as LLM has good performance in the test set or the real world. Feldman (2020) has conducted a theoretical analysis explaining why memorization happened and is needed. They showed in a long-tail distribution where many categories only have very few data. In this situation, as the number of data does not support extracting general features to describe the long-tail data, the best choice for the neural model is just to memorize them and compare them with data in the test set.
In the era of LLM, which differs from previous classification models, the LLM can directly output their pre-training data also known as memorization, previous researchers extracted knowledge bases from LLM Petroni et al. (2019); AlKhamissi et al. (2022) showing massive knowledge memorized by LLM. However, the massive knowledge learned during pre-training also leads to the concern of data contamination Sainz et al. (2023) and privacy risk Yao et al. (2024). Regarding specific research of memorization behavior itself, Carlini et al. (2023) discussed memorization behavior in LLM regarding their factors of model size, continuation size, and context size, suggesting they are inter-correlated. Biderman et al. (2023a) also discussed memorization in LLM regarding their training phase and the overlap between different model sizes, showing the possibility of using smaller LLM to predict memorization of larger LLM.
Compared to the previous memorization studies, our research discussed memorized text and extended this scope to less and unmemorized text. Additionally, we also discussed both the encoding and decoding dynamics of generations. Finally, we analyze the prediction results of memorization.
Figure 1: Memorization and Research Scope
3 Experiment Seeting
3.1 Criteria
3.1.1 Memorization Criteria
We choose the K-extractability Carlini et al. (2021) to define the memorization and calculate the memorization score based on it.
In our setting, we prompt the model with a specific context length and use greedy decoding to generate the continuation tokens based on that previous context. Then, we compare the predicted continuations with the actual continuations and calculate a memorization score using the following equation:
$M(X,Y)=\dfrac{\sum_{i=0}^{n}x_{i}==y_{i}}{L(Y)}$ (1)
In the above equation, $X=\{x_{0}\cdots x_{n}\}$ is the sequence of predicted continuation tokens, and $Y=\{y_{0}\cdots y_{n}\}$ is the sequence of true continuation tokens. $L(Y)$ means the length of the true continuation tokens. This score is the memorization score of the sequence $Y$ under context tokens $C=\{c_{0}\cdots c_{n}\}$.
From this equation, we can see that if $M(X,Y)==1$ means that the sequence $Y$ is completely memorized under the context sequence $C$, which we call this sentence K-extractable. Similarly, we can also say that the sequence $Y$ is completely not memorzied if $M(X,y)==0$ In addition to this token level-based accuracy, we collected all sentences and divided them into different ranges in our experiment.
3.1.2 Prediction Criteria
In this study, we also perform the prediction of memorization. We use two criteria to evaluate prediction performance, e.g., Token-Level Accuracy and Full Accuracy.
We define the prediction at a specific token at idx $i$ as $\hat{X}=\{\hat{x_{0}},\hat{x_{1}}\dots\hat{x_{n}}\}$ where each prediction $\hat{x_{n}}$ is a binary label indicating the token at this index is a memorized token or not. Similarly, the gold result is denoted as $\hat{Y}=\{\hat{y_{0}},\hat{y_{0}}\dots\hat{y_{n}}\}$ where is $\hat{y}$ is the golden binary label
The Token Level Accuracy can be defined as:
$T(\hat{X},\hat{Y})=\dfrac{\sum_{i=0}^{n}\hat{x_{i}}==\hat{y_{i}}}{L(\hat{Y})}$ (2)
The Full Accuracy means that the prediction of token level at a sentence must be entirely correct, which means $T(\hat{X},\hat{Y})$ must equal $1$.
3.2 Model Setting
We use the Pythia model Biderman et al. (2023b) to analyze the memorization behavior as Pythia provides LLMs trained across various sizes (70m to 12b) with the same training order using open-sourced Pile Gao et al. (2020) corpora. This provides the benefits of analyzing memorization behavior with a more stable model choice and experiment stability.
Specifically, we investigate the model size of [70m, 160m, 410m, 1b, 2.8b, 6.9b, 12b] where $m$ stands for million and $b$ stands for billion. We choose the model trained on the deduplicated pre-train data version to avoid the effect of duplicated sentences, as previous research reports that the chance to be memorized grows exponentially with the number of duplicates. . [^1] For more details, please refer to https://github.com/EleutherAI/pythia
4 Experiment Results
4.1 Memorization Factors
This section discusses how the number of memorized sentences grows with different factors. We collect the number of sentences memorized or unmemorized across different extend under different model sizes. The results are plotted in Figure 2.
Firstly, we can see that the memorized sentences increase with the model size and context size but decrease with the increase of complement size, which aligns with previous research Carlini et al. (2023). We illustrate a more in-depth analysis of those factors in the following sections.
4.1.1 The Factor of Model Size
In this experiment, we discuss how model size affects the number of memorized and unmemorized contexts. From the Figure 2 (a), we can see that
  • We can see that sentences with low memorization scores are significantly higher than sentences with high memorization scores. This shows that even though we can find memorization in the LLM, most of the contents they are trained in are not memorized.
  • Among sentences with different memorization scores, the memorized sentence increases faster than others. This indicates that LLM tends to memorize sentences more than to remember a piece of it fully. We can also see that the number of sentences with low memorization scores, like 0 and 0.1, decreases with the increase of model size, showing that the LLM tends to memorize the corpora with the increase of model.
  • Additionally, we can see that the sentences that are being memorized or memorized do not scale linearly with the increase in model size. For example, we can see that for fully memorized sentences, the increase in sentence numbers is obvious from 70m to 2.8b compared to the increase from 2.8b to 12b. Similarly, the same decreasing trend can be seen from the figure for the memorized sentence. This shows that even with enough model size, the LLM may still be unable to memorize the whole corpora.
Figure 2: Memorization Statistics Across Model Size, Complement Size, and Contenxt Size
4.1.2 The Factor of Context Size and Complement Size
We change the number of prompted tokens while fixing the complement size and then observe the memorization. From the Figure 2 (b) and (c), we can tell that:
  • The decrease in memorized sentences with the complement size is not linear. We can see a noticeable decrease when changing the complement size from 32 to 48, where the number of memorized sentences is almost halved.
  • Additionally, from this figure, we can see that some sentences are firmly memorized, and even when we increase the complement size from 64 to 96, the number of memorized sentences does not change too much.
  • We can also see that the drop in the number of memorized sentences is more significant in the larger model. This suggests that the increase in model size significantly helps to memorize short context compared to longer content.
  • The increase of memorized sentences with the context size is also not linear. The increase in context leads to an above-linear increase in the number of memorized sentences. This suggests that with more information on the context, the more LLM can generate related continuations.
Figure 3: Transition Across Different Model Size
4.2 Memorization Transition Across Model Size
This section discusses how different memorized sentences transit between different model sizes. For example, when trained with a larger model, we study how many unmemorized sentences in a small-size model will transition to memorized or low-memorized sentences. We plot the transition matrix in Figure 3. This figure shows the transition from the 410m size model to the 2.8b size and 2.8b size to the 12b size model. We assigned sentences with memorization scores in the range of 0.2 with labels of very low, low, medium, high, and very high. From the Figure 3, we can conclude the following:
  • Most sentences will stay at their previous state when trained in a larger model. We can tell this by looking at the diagonal of those two transition matrices. Most sentences will stay in their original states.
  • Additionally, highly memorized sentences are harder to transit compared to low memorized sentences. Especially for highly memorized sentences, more $94\%$ of sentences will remain being memorized compared to low memorized sentences or unmemorized sentences.
  • We can also see with the increase in the model size, the sentence is more inclined to stay in its original state. Figure b shows the transition from the 2.8b model to the 12b model; we can tell that the transition probability at the diagonal is higher than the 410b to 2.8b transition matrix counterpart. This suggests that the memorized or unmemorized sentences tend to be fixed with the increase in model size.
  • We can also see that even with a highly memorized sentence, there is still a little chance that the sentence will be forgotten or poorly memorized. The sentence in low memorized sentence tends to be highly memorized.
4.3 Input Dyanamics
In this section, we discuss the question of whether there is any sign when the model starts to generate memorized or unmemorized content. Especially what makes sentences memorized at different states different and why some sentences are memorized by large-size models but not small-size models.
4.3.1 What triggers the memorized or unmemorized content?
Firstly, we begin with the input level, which is the pre-train corpora. We conduct this study by analyzing the n-gram statistics. First, we show the dynamics of 1-gram statistics in the
Figure 4: One-gram Analysis at Each Index
From the statistics, we can see:
  • We can observe a very obvious boundary effect in this figure. At the index around 32 which is the first generated token, the frequency drops and rises for the memorized sentence (positive boundary effect) while rises and drops for the unmemorized sentences (negative boundary effect) while the frequency drop is more significant than the increase in the frequency of the memorized sentence.
  • For the half-memorized sentence, we see the frequency increase and drop around the index of 39 which is half the length of generated tokens. This shows that for the half-memorized sentence, the previous half is mostly memorized and the later half is mostly unmemorized meaning that the memorized tokens are distributed in a near continuous way rather than scatter in the generated tokens. We also notice the repetitive pattern of half-memorized sentences and unmemorized sentences which is that before the generation of the unmemorized token, we can see an increase and drop in frequency.
  • The positive boundary effect of a memorized sentence suggests that the memorized sentence is caused by the first several tokens having higher frequency showing the frequency of the initial tokens is more important than the following tokens, e.g remember the first several tokens make the following sentence to be memorized more easily. The negative boundary effect of the unmemorized sentences indicates the same results that the low frequency of the first several tokens makes the following sentence also easy to forget.
4.3.2 Model Size and the Role of Boundary Effect
We have shown the existence of a boundary effect in the above experiment. In this section, we discuss how such boundaries exist across model sizes and whether there are other factors influencing memorization. We calculate the average frequency of context, continuation, and also the frequency difference at the boundary in Table 1. From this table, we can obtain the following:
  • In the context and continuation, we can see the boundary effect even when we scale the calculation to the whole sentences. That is for memorized sentences, the frequency is higher in lower in the context but higher in the continuation while the unmemorized sentence shows the opposite.
  • However, the half-memorized sentence shows different results in that the frequency of continuation tokens is higher than the average frequency of context which is caused by the increase of frequency before reaching the first generated unmemorized tokens indicated in Figure 4.
  • In the specific boundary difference, we can see a decreasing trend for the positive boundary effect of memorized sentences while an increasing trend for the negative boundary effect of unmemorized sentences. These results suggest sentences with significant frequency increase are easy to memorize or forget.
Size Context Continuation Bounrady Freq Difference
One(B) Two(M) One(B) Two(M) One(B) Two(M)
M H U M H U M H U M H U M H U M H U
160m .708 .713 .744 55 69 .739 .837 .628 54 56 .12 -.22 -.78 3.3 -6.6
1b .713 .711 .752 55 69 .736 .832 .631 55 56 .11 -.27 -.76 2.8 -6.9
6.8b .721 .710 .759 .736 .829 .638 .09 -.15 -.74
12b .721 .720 .760 .736 .846 .626 .04 -.24 -.72
Table 1: Gram Statistics. $B$ means billion. $M$ means a million. We ignore the unit cardinal digit for the one-gram statistics as the unit digit is the same, which is $1$ for Context and Continuation while 00 for Freq Diff.
4.4 Output Dynamics
4.4.1 Embedding Dynamics
In this section, we discuss the embedding of generated tokens for sentences memorized to different extents. We collect the hidden state at the last layer of each token. We compute the pair-wise Euclidean distance and Cosine similarity between sentences memorized at different extents across each generation step. Then, we perform dimension reduction on the embedding and visualize it in the 2D space using PCA.
Figure 5: Embedding Dynamics Across Different Model Size. Score $x$ means $x$ tokens are memorized.
From the Figure 5, we can obtain the following points:
  • The cosine similarity remains relatively stable across different steps between sentences memorized at different extents. This shows that different memorized sentences have different parts in the embedding space and remain relatively stable.
  • Meanwhile, we can see the Euclidean distance decreases with the generation of tokens. Combing with the previous observation that sentences memorized with different extend have a stable Cosine Similarity, we can conclude that with the generation of tokens, the angle between sentence vectors at high dimension space remains stable while the cosine distance decreases, which suggests the sentence vectors is shortened while the angle remains stable.
  • Additionally, we can see that the large-size model has a higher Euclidean distance and lower Consine Similarity. The high Euclidean distance is caused by the expansion of hidden size (512 for 70m, 2048 for 1b, etc.). As for the decrease of low Consine Similarity, this can also be attributed to the hidden size as the increase in hidden size expands the embedding space. This gives more expressivity to the embedding space, as the Euclidean distance between sentence embedding also expands, which leads to the decrease of Cosine Similarity. This may also help to explain the performance gap between different model sizes. Large-size models have more expressivity in the embedding space; thus, different sentences with different meanings are more clearly distributed with fewer neighbors in the space. The smaller embedding space in a small model makes the mix of embedding easier, leading to ambiguity in the embedding space.
4.4.2 Generation Dynamics and Entropy
In this section, we discuss how the model generates the memorized and unmemorized sentences, e.g., the generation dynamics. This means that the model exhibits certain behaviors when it generates memorized sentences or unmemorized sentences. Or does the LLM generate memorized and unmemorized sentences in the same way? To discuss this problem, we collect the generation probability at each token when the LLM generates memorized, unmemorized, and half-memorized sentences, calculate the entropy across 10,000 sentences, and plot the entropy at each token in Figure.
Figure 6: Averaged Entropy at Each Index
From the Figure 6 we can tell:
  • First, we can see that the entropy is different depending on the extent of memorization. We can clearly see that an unmemorized sentence has a higher average entropy at each token than an unmemorized sentence. This shows that when the LLM generates unmemorized sentences, they are more unconfident about generating tokens.
  • Additionally, we can tell that when the model starts to generate unmemorized tokens, we can see a significant increase in entropy. This boundary effect is very obvious. This suggests that when the model generates unmemorized sentences, their entropy will suddenly increase instead of decrease.
  • We can also notice that the boundary for generating unmemorized sentences is mostly consistent, which means we did not see multiple boundaries. This suggests that even for half-memorized sentences, the model tends to generate continuous memorized tokens then, followed by unmemorized tokens instead of generating mixed memorized and unmemorized tokens. This suggests that the LLM memorizes sentences at a chunk level rather than a word level.
  • We can also see that the entropy will decrease with the increase in model size. This suggests that the larger size model is more confident about the generation of tokens compared to the small size model.
Length 70m 410m 1b 2.8b 6.9b 12b
Token Full Token Full Token Full Token Full Token Full Acc Token
16 78.2 10.2 78.6 10.4 78.8 10.6 80.1 10.7 77.4 8.3 80.3 10.9
32 78.6 5.9 79.6 6.0 79.7 6.1 80.1 6.3 80.5 6.4 80.8 6.4
48 79.6 5.2 80.3 5.4 80.4 5.6 80.4 5.5 80.8 5.8 81.0 6.0
64 80.1 4.7 80.8 4.8 81.2 5.2 81.5 5.5 81.8 5.8 82.1 6.0
Table 2: Performance of Transformer Model on Prediction of Memorization. The context length is defaulted at 32. Length means the token length that is required to be predicted. Token and Full represent Token-Level Accuracy and Full Accuracy. The higher, the better. The best results across model sizes are bold, and the best results within a model size are underlined.
4.5 Prediction of Memorization
In this section, we train a Transformer model to predict the memorization of tokens from the model’s embedding and statistic features and analyze the results. We discuss whether Transformer is capable of this task and the difficulty of predicting the memorized token at different model sizes as in the real world, similar context sentences can also trigger memorized texts Stoehr et al. (2024) and the generated token can also be a paraphrase of the actual continuation Ippolito et al. (2023).
4.5.1 Results on Memorization of Prediction
The training is formalized by receiving both the last layer’s embeddings at each step and statistics (e.g., the entropy) then the Transformer is required to generate a binary classification label at each index, predicting whether the token at this index is memorized. The results are presented in Table 2. From the table, we can obtain the following results:
  • First, regarding Token-level accuracy, we can see with a naive Transformer model, the token-level accuracy can reach $80\%$ and even higher, showing that the memorization label of $80\%$ tokens can be correctly classified. The Full-level accuracy is relatively low as this requires a correct prediction at every token. However, the high token level accuracy at such a naive model shows the potential that prediction of memorization at the decoding step is possible.
  • Secondly, we showed that with the increase of the model size of LLM, the Token-level accuracy and Full-level accuracy increase rather than drop. This suggests that it is easier to predict the memorization behavior of larger-size models rather than small-size models.
  • Thirdly, we found that with the increase in continuation size, the Token level accuracy gradually increases, possibly due to the increase in training data. For example, if the length to be predicted is $64$, the token used to train the model is four times higher than the prediction length of $16$. However, we see a drop in the Full accuracy as the increase in prediction length leads to more tokens to be predicted; thus, reaching Full accuracy is harder in a longer prediction length.
4.5.2 Analysis of Full Accuracy
Figure 7: Distribution Across Model Szize of Full Accurate Predictions
In this experiment, we use the trained Transformer model to generate a case study for different memorized sentences. We presented the predicted ability for the corresponding token to be memorized. A higher probability means the token here is memorized, and a lower probability means the token here is unmemorized. The case study is presented in Figure 8. From the figure, we can tell:
  • For any model size, most sentences that are fully predicted share low memorization scores. Especially, the 00 score (unmemorized) sentence occupies the largest portion among fully predicted sentences. Meanwhile, we can see that sentences with high memorization scores are hard to be fully predicted. We can see that the $1$ score (memorized) sentence has the lowest share among fully predicted sentences, showing that memorized sentences are hard to predict.
  • Across model size, we can see that with the increase in model size, the proportion for unmemorized sentences increases. This suggests that the prediction model in the larger model is more accurate in predicting unmemorized sentences rather than memorized sentences.
  • From previous Figure 6, for the memorized sentence, the entropy decrease after decoding is less significant than for the unmemorized sentence, whose entropy will suddenly increase. Additionally, the decrease in entropy is less significant with the increase in model size as the large model is more confident about the decoding, leaving less space for the entropy decrease. Since entropy is part of the feature we input to the model, the aforementioned entropy decreases, making a prediction of memorized sentences easier. The less significant decrease of entropy in memorized sentences helps to explain why the prediction of memorized sentence
4.6 Conclusion
In this research, we comprehensively discussed the memorization behavior of LLMs across different perspectives. From a general level, we conclude statistics regarding the memorization behavior of LLMs, which not only aligns with previous research but also extends to sentences with lower memorization scores. From the embedding level, we obtained the embedding dynamics of sentences with different memorization scores. We found that the cosine similarity stays stable among sentences with different memorization scores, but the Euclidean distance decreases. The increase in model size leads to a decrease in cosine similarity and an increase in Euclidean distance. At the decoding dynamics, we conducted an analysis at the distribution level and calculated the entropy. We found that when the LLM starts to generate an unmemorized sentence, the entropy will suddenly rise at the boundary between memorized and unmemorized parts. On the contrary, when the model generates memorized sentences, the entropy will also have a sudden drop, though not obvious, comparing the generation of unmemorized parts. Finally, we trained a Transformer model to predict the memorization. We found that, Through a detailed case study, we analyzed the prediction behavior and showed that context-based memorization prediction is possible.
5 Limitations
This research has analyzed various factors regarding the memorization phenomena of LLMs. We acknowledge that there are still limitations to this research. Due to the lack of LLMs whose models and data are both being released, as the data is also different, it is hard to compare the memorization across different LLMs. How to comprehensively measure memorization of various LLMs, either close-sourced like GPT-4 or open-sourced like Pythia or LLaMa, can be a future research. Additionally, Pythia only provides LLMs up to 12B, which is still smaller than SotA open-sourced LLMs like LLaMA, whose largest model is 70B. The difference between 12B and 70B may lead to emergent abilities, which may also affect the memorization results obtained in 12B, which may not be able to scale even further due to those factors. Though we trained the Transformer to predict the memorization of LLMs, it is more analysis-oriented; thus, the performance is not the main focus of this experiment. Additionally, in this research, the discussion of memorization is under verbatim memorization, where generated tokens are identical to the same sentence in the corpora.
Appendix A Example Appendix
Figure 8: Prediction Examples
This is an appendix.