### v3

# A Multi-Perspective Analysis of Memorization in Large Language Models

Abstract
The memorization behavior of Large Language Models (LLMs) means that LLMs can generate the same content used to train them. Though previous research has discussed memorization, the memorization of LLMs still lacks explanation, especially the cause of memorization and the dynamics of generating them. In this study, we comprehensively discussed memorization from various perspectives and extended the discussion to less and unmemorized content. Through various studies, we found that: (1) Through experiments, we revealed the relation of memorization between model size, continuation size, and context size. Further, we showed the dynamics of how unmemorized sentences transition to memorized sentences. (2) Through input analysis, we found the frequency of tokens related to memorization and discovered the boundary effect when generating memorized or unmemorized sequences. (3) Through an analysis of embedding and decoding, we showed the LLM can generate paraphrased memorized sentences and sentences of different memorization scores clustered in the embedding space. The entropy also shows a boundary effect when generating memorized or unmemorized content. (4)We trained a Transformer model to predict the memorization of different models, showing the possibility of predicting memorization and paraphrase memorization. Further analysis shows that the prediction of unmemorized sequences is easier than the prediction of memorized content, which can be explained by the significance of the boundary effect.
1 Introduction
Large language models (LLMs), spanning from BERT Devlin et al. (2019) to the recent strongest GPT-4 OpenAI et al. (2024), have not just revolutionized the Natural Language Processing (NLP) but nearly the whole artificial intelligence research field. Though their performances on various tasks are surprising, we still know little about the underlying mechanism of Large Language Models whose structures are based on neural models known as black-box Alain and Bengio (2018). Due to the enormous model size and pre-train data size, unprecedented phenomena in previous machine learning are also identified in the LLM, while in this research, we discuss the memorization.
Memorization Hartmann et al. (2023) in the context of LLMs means the LLM sometimes will verbatim output the same sentence recorded in their pre-training corpus. Memorization is a coin with two sides. On the one hand, utilizing such memorization, we can use LLMs as a knowledge base Petroni et al. (2019) that can be served for multiple uses. On the other hand, personal information contained in the pre-train corpora may also be elicited maliciously Yao et al. (2024). Though such a phenomenon has been utilized or studied since the merge of LLMs, detailed analysis of memorization is still under-explored.
In this research, we discuss memorization in detail and analyze it from several perspectives, e.g., the pre-train data level, the encoding level, the generation level, and finally, we try to predict memorization. Following those studies, we found that:
  • For both memorization and unmemorized sentences, even though their numbers increase/decrease with the model size, the trend is not linear, showing the maximum capacity to memorize sentences exists. The number of memorized sentences decreases sub-linearly with the continuation size while increasing above linearly with the context size. We also study how unmemorized sentences transition to memorized sentences and vice versa. We found that memorized sentences are more likely to be memorized when scaling the model size, while low-memorized sentences will be gradually memorized.
  • Through a study of the encoding dynamics, we found that sentences with different memorization scores are clustered in the embedding space. Additionally, with the generation of tokens, the embedding similarly remains basically stable while the Euclidean distance is shrinking. Moreover, the larger the model, the more significant the difference in the Cosine Similarity. This may help to explain why the larger model performs better than the smaller-size model.
  • In both the n-gram and decoding entropy dynamics, we found that a boundary phenomenon exists when the model generates memorized or unmemorized sentences. In the n-gram, the frequency when generating unmemorized tokens decreases while increasing for memorized tokens. In the entropy analysis, the entropy suddenly increases when unmemorized sentences are generated, while it decreases for memorized sentences.
  • Finally, we trained a naive Transformer model to predict memorization based on context. We show that prediction accuracy over tokens is high and grows with the model size, while it is still hard to predict correctly for a whole sentence
2 Related Works
2.1 Scaling Laws of LLM
In this study, our experiments span across various model sizes, thus relating to the research of Scaling Laws Kaplan et al. (2020); Abnar et al. (2021); Villalobos (2023), which suggests the LLM’s performance scales with the size of corpora, the parameter size and the computation required.
Those researches inspired LLMs to continuously scale those factors to gain higher performance like T5 Raffel et al. (2023), GPT-3 Brown et al. (2020), PaLM 2 Anil et al. (2023), etc. On the other hand, researchers also analyzed how LLM Scaling affects particular tasks rather than general performance. For example, Sun and Miceli-Barone (2024) discussed how LLM scaling affects both translation tasks and prompt injection attacks. Within those discussions, the LLM’s emergent abilities Wei et al. (2022) are also discovered, which means the LLM suddenly reaches high performance on previous low-performance tasks when the LLM reaches a specific model size such as digital addition Nye et al. (2021), reasoning Lampinen et al. (2022). Recent studies have questioned whether emergent abilities are just mirages caused by misuse of metrics Schaeffer et al. (2023) or whether emergent abilities are just context-learning Lu et al. (2023).
In the field of memorization, Carlini et al. (2023) discussed the memorization across model size and found that the number of memorized texts grows with the model size and context size. Additionally, Biderman et al. (2023a) also discussed similar topics and found that a large portion of memorized text in a small-size model is also memorized by a larger model, showing that memorized texts may share certain common features.
2.2 Memorization
Before LLM, a phenomenon close to memorization in real-world practices is known as over-fitting Tetko et al. (1995), which means a near-zero loss in the train set while bad performance in the test set. Zhang et al. (2017) described that the neural model could reach near-zero loss on the training set, which suggests the neural model perfectly memorized the relation between input and data label. However, memorization is different from over-fitting as LLM has good performance in the test set or the real world. Feldman (2020) has conducted a theoretical analysis explaining why memorization happens and is actually needed. They showed in a long-tail distribution where many categories only have very few samples. As the neural model cannot extract general features from the limited number of samples in those categories, the best choice for the neural model is just to memorize them and compare them with data in the test set.
In the era of LLM, different from classification models, the LLM can directly output their pre-training content that can be used from a knowledge graph Petroni et al. (2019); AlKhamissi et al. (2022) while also leads to the concern of data contamination Sainz et al. (2023) and privacy risk Yao et al. (2024). Regarding specific research on memorization behavior, Tirumala et al. (2022) discussed training dynamics of memorization, showing large models tend to memorize training samples more easily. Carlini et al. (2023) discussed memorization behavior in LLM regarding their factors of model size, continuation size, and context size, suggesting they are inter-correlated. Biderman et al. (2023a) also discussed memorization in LLM regarding their training phase and the overlap between different model sizes, showing the possibility of using smaller LLM to predict memorization of larger LLM.
Compared to the previous memorization studies, our research discussed memorized text and extended this scope to less and unmemorized text. Additionally, we also discussed both the encoding and decoding dynamics of generations. Finally, we trained a Transformer model to predict memorization and analyze the results.
Figure 1: Memorization and Research Scope
3 Experiment Seeting
3.1 Criteria
3.1.1 Memorization Criteria
We choose the K-extractability Carlini et al. (2021) to define the memorization and calculate the memorization score based on it.
We prompt the LLM with a sequence of context tokens $C=\{c_{0}\cdots c_{n}\}$ whose length is defined as context length and use greedy decoding to generate the continuation tokens based on the prompted context. Then, we compare the predicted continuations with the actual continuations and calculate a memorization score using the following equation:
$M(X,Y)=\dfrac{\sum_{i=0}^{n}x_{i}==y_{i}}{L(Y)}$ (1)
In the above equation, $X=\{x_{0}\cdots x_{n}\}$ is the sequence of predicted continuation tokens, and $Y=\{y_{0}\cdots y_{n}\}$ is the true continuation tokens. $L(Y)$ means the length of the true continuation tokens. If $M(X,Y)==1$, this sequence is completely memorized under the context sequence, which is called K-extractable. Similarly, we also define the sequence $Y$ as completely not memorized if $M(X,y)==0$. Since the memorization score has high granularity, in some experiments, we also classify sentences into several memorization levels based on their memorization scores.
3.1.2 Prediction Criteria
In this study, we use two criteria to evaluate model performance in predicting memorization, e.g., Token-Level Accuracy and Full Accuracy.
We define the prediction for a sequence of generated tokens as $\hat{X}=\{\hat{x_{0}},\hat{x_{1}}\dots\hat{x_{n}}\}$ where each prediction $\hat{x_{n}}$ is a binary label indicating the token at this index is a memorized token or not. Similarly, the gold result is denoted as $\hat{Y}=\{\hat{y_{0}},\hat{y_{0}}\dots\hat{y_{n}}\}$ where is $\hat{y}$ is the golden binary label.
The Token Level Accuracy can be defined as:
$T(\hat{X},\hat{Y})=\dfrac{\sum_{i=0}^{n}\hat{x_{i}}==\hat{y_{i}}}{L(\hat{Y})}$ (2)
The Full Accuracy means that the prediction of token level at a sentence must be entirely correct, which means $T(\hat{X},\hat{Y})$ must equal $1$.
3.2 Model Setting
We use the Pythia model Biderman et al. (2023b) to analyze the memorization behavior as Pythia provides LLMs trained across various sizes with the same training order using open-sourced Pile Gao et al. (2020) corpora, which benefits the experiment and provides stability.
Specifically, we investigate the model size of [70m, 160m, 410m, 1b, 2.8b, 6.9b, 12b] where $m$ stands for million and $b$ stands for billion. We choose the model trained on the deduplicated pre-train data version to avoid the effect of duplicated sentences, as previous research reports that the chance to be memorized grows exponentially with the number of duplicates Kandpal et al. (2022). [^1] For more details, please refer to https://github.com/EleutherAI/pythia
4 Experiment Results
4.1 Memorization Factors
This section discusses how memorization is affected by different factors. We collect the number of sentences memorized across different extend under different model sizes. Then, we divide them into ten sets with a memorization difference of 0.1 The results are plotted in Figure 2.
Firstly, we can see that the memorized sentences increase with the model size and context size but decrease with the increase of complement size, which aligns with previous research Carlini et al. (2023). We illustrate a more in-depth analysis of those factors in the following sections.
4.1.1 The Factor of Model Size
In this experiment, we discuss how model size affects the number of memorized and unmemorized contexts. From the Figure 2 (a), we can obtain that:
  • We can see that the number of sentences with low memorization scores (0-0.3) is significantly higher than sentences with high memorization scores. This shows that though memorization exists in the LLM, most of the pre-train data are still not memorized.
  • Among sentences with high memorization scores, the fully memorized sentence increases faster than others. This indicates that LLM tends to fully memorize sentences rather than memorizing a piece of it. We can also see that the number of sentences with low memorization scores (0, 0.1, etc) decreases with the increase in model size. This shows that unmemorized sentences will gradually memorized with the increase in model size.
  • Additionally, we can see that the sentences that are being memorized or memorized do not increase or decrease linearly with the increase in model size. For fully memorized sentences, the increase in sentence numbers is obvious from 70m to 2.8b compared to the increase from 2.8b to 12b. Similarly, the same decreasing trend can be seen from this figure for the unmemorized sentence. This shows an upper limit exists for memorization that the LLM cannot memorize the whole corpora even with enough model size.
Figure 2: Memorization Statistics Across Model Size, Complement Size, and Contenxt Size
4.1.2 The Factor of Context Size and Complement Size
We change the number of prompted tokens while fixing the complement size and then observe the memorization. From the Figure 2 (b) and (c), we can tell that:
  • The decrease in memorized sentences with the complement size is not linear. When we increase the complement size from 64 to 96, the number of memorized sentences does not change too much compared to the change when the complement size changes from 32 to 48. This also suggests some sentences are firmly memorized.
  • Additionally, the drop in the number of memorized sentences with the increased complement size is more significant in the larger model. This shows that though larger models memorize more sentences, the memorized sentence is not as firm as sentences memorized by small models.
  • The increase of memorized sentences with the context size is also not linear. The increase in context leads to a near-exponential increase in the number of memorized sentences. Additionally, the increase is more significant in the larger model, showing that more potential memorized contents can be elicited by giving more context compared to the small model.
Figure 3: Transition Across Different Model Size
4.2 Memorization Transition Across Model Size
This section discusses how sentences with different memorization scores transit across model sizes. Specifically, when trained with a larger model, we study the mutual transition between sentences with different memorization scores. We classify sentences with memorization scores with a range difference of 0.2 with labels of very low, low, medium, high, and very high. We plot the transition matrix in Figure 3, which shows the transition from the 410m size model to the 2.8b size and 2.8b size to the 12b size model.
From the Figure 3, we can conclude:
  • Most sentences will stay at their previous state even when trained in a larger model. We can tell this by looking at the diagonal of those two transition matrices. Additionally, the higher the memorization score, the harder the sentence will transition to another state. For highly memorized sentences, more than $90\%$ of sentences will remain being memorized compared to low memorized sentences.
  • We can also see with the increase in the model size, the sentence is more inclined to stay in its original state. In the transition from the 2.8b model to the 12b model, we can tell that the transition probability at the diagonal is higher than the 410m to 2.8b transition matrix counterpart. This suggests that the memorized or unmemorized sentences tend to be fixed with the increase in model size.
  • We also see that even for a highly memorized sentence, there is still a little chance that the sentence will transit to a low memorized state, suggesting some sentences may be memorized randomly rather than following a specific feature. Such randomness makes it possible for them to transit to a low memorized state when trained in a large model.
4.3 Input Dyanamics
In this section, we discuss the question of whether there is any sign when the model starts to generate memorized or unmemorized content. Especially what makes sentences memorized at different extents and why some sentences are memorized by large-size models but not small-size models.
4.3.1 Token Level Frequency Analysis
Figure 4: One-gram Analysis at Each Index
Firstly, we begin with the input level by analyzing the n-gram statistics of the pre-train corpora. We show the 1-gram statistics across steps in Figure 4, and we can see:
  • We can observe a very obvious boundary effect. At the index around 32, which is the first generated token, the frequency drops and rises for the memorized sentence (positive boundary effect) while rises and drops for the unmemorized sentences (negative boundary effect). Additionally, the negative boundary effect is more obvious than the positive boundary effect.
  • For the half-memorized sentence, we see the frequency increase and drop around the index of 39, which is half the length of generated tokens. This shows that for the half-memorized sentence, the previous half is mostly memorized, and the later half is mostly unmemorized, meaning that the memorized tokens are distributed in a near continuous way rather than scattered in the generated tokens. We also notice the repetitive pattern of half-memorized sentences and unmemorized sentences, which is the negative boundary effect around the index of 39.
  • The positive boundary effect of a memorized sentence suggests that the memorized sentence is caused by the first several tokens having a higher frequency, showing the frequency of the initial tokens is more important than the following tokens, e.g., remember the first several tokens make the following sentence to be memorized more easily. The negative boundary effect of the unmemorized sentences indicates the same results: the low frequency of the first several tokens makes the following sentence easy to forget.
4.3.2 Sentence Level Frequency Analysis
We have shown the existence of a boundary effect in the above experiment. In this section, we discuss how such boundaries exist across model sizes and whether there are other factors influencing memorization. We calculate the average frequency of context, continuation, and also the frequency difference at the boundary in Table 1. From this table, we can obtain the following:
  • First, we can see that the frequency of one-gram is significantly higher than the frequency of two-gram. In the context and continuation, the average frequency is roughly 3.5 times higher than the two-gram frequency. Additionally, such boundary effect is consistent in the two-gram setting. However, the positive boundary effect is less obvious due to the decrease of frequency when computed by 2-gram. However, as the unit is billion, the actual frequency gap is still a million level.
  • In the context and continuation, in memorized sentences, the frequency is higher in lower in the context but higher in the continuation, while the unmemorized sentence shows the opposite. This suggests the boundary effect also exists in a broader sequence level, though less obvious.
  • However, the half-memorized sentence shows different results in that the frequency of continuation tokens is higher than the average frequency of context, which is caused by the increase of frequency before reaching the first generated unmemorized tokens indicated in Figure 4.
  • In the specific boundary frequency difference around the first generated token, we can see a decreasing trend for both positive and negative boundary effects. However, the decrease in the positive boundary effect makes it less significant, while the decrease in the negative boundary effect makes it more significant. These results suggest that the significance of the positive boundary corresponds to how easily the sentence can be memorized. In contrast, the significance of the negative boundary effect corresponds to how easily the sentence can not be memorized.
Size Context Continuation Bounrady Freq Difference
One Two One(B) Two One Two
M H U M H U M H U M H U M H U M H U
160m 1.708 1.713 1.744 0.551 0.534 0.691 1.739 1.837 1.628 0.535 0.659 0.567 0.114 0.330 -0.939 0.033 0.101 -0.663
1b 1.713 1.711 1.752 0.558 0.552 0.697 1.736 1.832 1.631 0.509 0.682 0.564 0.103 0.270 -0.981 0.028 0.090 -0.696
6.8b 1.721 1.710 1.759 0.570 0.565 0.701 1.736 1.829 1.638 0.496 0.699 0.564 0.090 0.140 -0.963 0.027 0.085 -0.726
12b 1.721 1.720 1.760 0.572 0.569 0.702 1.736 1.846 1.626 0.493 0.704 0.563 0.039 0.237 -1.016 0.026 0.083 -0.732
Table 1: Gram Statistics. The frequency unit is billion. Boundary Freq Difference means we use the frequency of the generated token to subtract the frequency of the last token in the context.
4.4 Output Dynamics
4.4.1 Embedding Dynamics
In this section, we discuss the embedding of generated tokens for sentences memorized at different extents. We collect the hidden state at the last layer of each token. We compute the pair-wise Euclidean distance and Cosine similarity between sentences memorized at different extents across each generation step. Then, we perform dimension reduction on the embedding and visualize it in the 2D space using PCA.
Figure 5: Embedding Dynamics of Across Different Model Size. Memorized Token Num $x$ means $x$ tokens generated to align with the true continuation.
From Figure 5, we can obtain the following points [^2] Detailed numbers of Cosine Similarity and Euclidean distance between sentences with different memorization scores are at different decoding steps are at Appendix B.2:
  • The cosine similarity remains relatively stable across different encoding steps between sentences memorized at different extents. Meanwhile, from the movement of those points, we can see the Euclidean distance decreases with the generation of tokens. Combined with the previous observation that sentences memorized with different extend have a stable cosine similarity, we can conclude that with the generation of tokens, the angle between sentence vectors at high dimension space remains stable while the cosine similarity decreases.
  • Additionally, we can see that for highly memorized sentences, even when they are not totally memorized. They are still close in the embedding space, indicating that the generated content may be both semantically and lexically similar to the memorized content.
  • Additionally, we can see that the large-size model has a higher Euclidean distance and lower Consine Similarity. The high Euclidean distance is caused by the expansion of hidden size (512 for 70m, 2048 for 1b, etc.). The decrease of the cosine similarity can also be attributed to the hidden size, as the increase in hidden size expands the embedding space. This gives more expressivity to the embedding space, as the Euclidean distance between sentence embedding also expands, which leads to the decrease of Cosine Similarity. This may also help to explain the performance gap between different model sizes. Large-size models have more expressivity in the embedding space; thus, different sentences with different meanings are more clearly distributed with fewer neighbors in the space. The smaller embedding space in a small model makes the mix of embedding easier, leading to ambiguity in the embedding space, thus degrading the performance.
4.4.2 Generation Dynamics and Entropy
Figure 6: Averaged Entropy at Each Index
In this section, we discuss how the model generates the memorized sentences with different memorization scores, e.g., the generation dynamics. As we have found the boundary effect in the input level, we also want to discuss the question of does the model exhibit certain behaviors when it generates memorized sentences or unmemorized sentences? To discuss this problem, we collect the generation probability at each token when the LLM generates memorized, unmemorized, and half-memorized sentences, calculate the entropy across 10,000 sentences, and plot the entropy at each token in Figure 6 and we can obtain the following from it:
  • First, we can see that the entropy is different depending on the extent of memorization. We can clearly see that an unmemorized sentence has a higher average entropy at each token than an unmemorized sentence. This shows that when the LLM generates unmemorized sentences, they are more unconfident about generating them.
  • Additionally, when the model starts to generate unmemorized tokens, we can see a significant increase in entropy, showing such boundary effect also exists in the decoding steps while behaving oppositely compared to the boundary effect in frequency analysis. We can also see a drop in entropy when decoding the memorized sequences. This corresponds to the fully memorized tokens and the first half of the half-memorized sequences. The entropy drops around the index of 32, which is the first generated token. However, we can see the drop of entropy for the memorized sentence is not samely significant as the half-memorized sentence, showing that even for the memorized content, the model is more confident when the sentence is fully memorized, leading to lower entropy.
  • We can also notice there is only one boundary at unmemorized or memorized sentences, further confirming that the model tends to memorize tokens continuously.
  • We can also see the entropy decrease with the increase in model size. This suggests that larger models are more confident about generating tokens than small ones. Additionally, we can see that the entropy is lower in the context of memorized tokens, showing that the entropy of context also relates to how a sentence is memorized.
Length 70m 410m 1b 2.8b 6.9b 12b
Token Full Token Full Token Full Token Full Token Full Acc Token
16 78.2 10.2 78.6 10.4 78.8 10.6 80.1 10.7 77.4 8.3 80.3 10.9
32 78.6 5.9 79.6 6.0 79.7 6.1 80.1 6.3 80.5 6.4 80.8 6.4
48 79.6 5.2 80.3 5.4 80.4 5.6 80.4 5.5 80.8 5.8 81.0 6.0
64 80.1 4.7 80.8 4.8 81.2 5.2 81.5 5.5 81.8 5.8 82.1 6.0
Table 2: Performance of Transformer Model on Prediction of Memorization. The context length is defaulted at 32. Length means the token length that is required to be predicted. Token and Full represent Token-Level Accuracy and Full Accuracy. The higher, the better. The best results across model sizes are bold, and the best results within a model size are underlined.
4.5 Prediction of Memorization
In this section, we train a Transformer model to predict the memorization of tokens from the LLM’s embedding and statistic features and analyze the results. We discuss whether Transformer can do this task and the difficulty of predicting the memorized token at different model sizes. As in the real world and indicated by our experiment results in previous sections, similar context sentences can also trigger memorized texts Stoehr et al. (2024), and the generated token can also be a paraphrase of the actual continuation Ippolito et al. (2023).
4.5.1 Results on Memorization of Prediction
The training is formalized by receiving both the last layer’s embeddings at each step and statistics (e.g., the entropy). Then, the Transformer is required to generate a binary classification label at each index, predicting whether the token at this index is memorized. The results are presented in Table 2. From the table, we can obtain the following results:
  • First, regarding Token-level accuracy, we can see with a naive Transformer model, the token-level accuracy can reach $80\%$ and even higher, showing that the prediction of memorization at the token level is easy. The full accuracy is low as this requires a correct prediction for every token.
  • Secondly, we showed that with the increase of the model size of LLM, the Token-level accuracy and Full-level accuracy increase rather than drop. This suggests that it is easier to predict the memorization behavior of larger-size models rather than small-size models as the embedding distance of large models is further, thus making the classification easier.
  • Thirdly, we found that with the increase in continuation size, the Token level accuracy gradually increases, possibly due to the increase in training data. For example, if the length to be predicted is $64$, the token used to train the model is four times higher than the prediction length of $16$. However, we see a drop in the Full accuracy as the increase in prediction length leads to more tokens being predicted. This makes reaching full accuracy harder as the tokens required to be predicted grow.
4.5.2 Analysis of Full Accuracy
Figure 7: Distribution Across Model Szize of Full Accurate Predictions
In this experiment, we use the trained Transformer model to generate a case study for different memorized sentences. We presented the predicted ability for the corresponding token to be memorized. A higher probability means the token here is memorized, and a lower probability means the token here is unmemorized. The case study is presented in Figure 8. From the figure, we can tell:
  • For any model size, most sentences that are fully predicted share low memorization scores. The 00 score (unmemorized) sentence occupies the largest portion of fully predicted sentences. Meanwhile, we can see that sentences with high memorization scores are hard to predict fully and correctly. We can see that the $1$ score (memorized) sentence has the lowest share among fully predicted sentences, showing that memorized sentences are hard to predict.
  • Across model size, we can see that with the increase in model size, the proportion for unmemorized sentences increases. This suggests that the prediction model in the larger model is more accurate in predicting unmemorized sentences rather than memorized sentences.
  • The reason why a sentence with a low memorization score can be easily predicted can be due to the difference of significance at the boundary. As obtained from our previous experiments, the boundary effect of unmemorized sequences is more significant than memorized sequences in either token frequency or entropy of the generated token. Especially we can see that the boundary effect of memorized sequences decreases with model size while increasing for unmemorized sequences in previous results. This helps to explain why predicting memorization is easier in small-size models but harder in large-size models because the boundary effect becomes insignificant when we increase the model size. Similarly, the increase of boundary effect for unmemorized sentences with model size makes them easier to predict.
4.6 Conclusion
In this research, we comprehensively discussed the memorization behavior of LLMs across different perspectives. From a statistical level, we analyzed the memorization behavior of LLMs and expanded the analysis to sentences with lower memorization scores. Further transition analysis revealed the transition dynamics when trained in a large model, showing that most sentences stay in the original state while the memorized state is hard to transition to the unmemorized state. At the input level, through frequency analysis, we found the positive and negative boundary effects when generating memorized and unmemorized tokens. We showed the LLM tends to memorize longer sequences rather than short pieces. In the output decoding level, in the embedding level, we found clusters of sentences with different memorization scores in the embedding space, and the close distance of sentences with high memorization scores indicates the existence of paraphrase memorization. We analyzed the distribution level and calculated the entropy at the decoding dynamics. We found that when the LLM starts to generate an unmemorized sentence, the entropy boundary effect also exists and significantly changes with the model size and the memorization score of the corresponding sentence. Finally, we trained a Transformer model to predict the memorization. We found that prediction at the token level is easy, while prediction at the sentence level is challenging. Through analysis of correctly fully predicted samples, we found unmemorized tokens are easy to predict while memorized tokens are hard to predict, and such a phenomenon can be explained by the significance of the boundary effect.
5 Limitations
This research has analyzed various factors regarding the memorization phenomena of LLMs. We acknowledge that there are still limitations to this research. Due to the lack of LLMs whose models and data are both being released, as the data is also different, it is hard to compare the memorization across different LLMs. Future research can be on how to comprehensively measure memorization of various LLMs, either close-sourced like GPT-4 or open-sourced like Pythia or LLaMa. Additionally, Pythia only provides LLMs up to 12B, which is still smaller than SotA open-sourced LLMs like LLaMA, whose largest model is 70B. The difference between 12B and 70B may lead to emergent abilities, which may also affect the memorization results obtained in 12B, which may not be able to scale even further due to those factors. Though we trained the Transformer to predict the memorization of LLMs, it is more analysis-oriented; thus, the performance is not the main focus of this experiment. Additionally, in this research, the discussion of memorization is under verbatim memorization, where generated tokens are identical to the same sentence in the corpora.
Appendix A Appendix
Figure 8: Prediction Examples. Pred Prob means the output predicted probability of the corresponding label in the pred row for each example. Gold means the true label, Pred means the predicted label, and Pred Prob means the probability of the corresponding prediction. $M$ means the label for a token at this index is a memorized token. $U$ means the label for a token at this index is an unmemorized token.
Appendix B Experiment Setting
The experiment uses 64 A100 40Gbs GPUs when using LLMs to generate tokens given the previous context, which utilizes PyTorch’s parallel running packages. We run the model with half-precision, which increases both the speed and saves memory. This follows the previous Pythia implementation when generating tokens given context.
The running time depends on the model size and the generated token length. In the situation when a 70m model with 32 context tokens and 16 tokens is required to be generated, it can be run with one A100 GPU within several hours. However, if such an experiment is in a single A100 GPU, it would estimated to take two weeks to finish the generation. Therefore, with 64 GPUs, the running of the 12b model in such a situation can be shortened to around one day. However, the generation time also largely increases with the length of generated tokens, which grows in a linear way with the number of tokens to be generated. Additionally, since we use greedy decoding and do not consider other possible decoding options but the token with the highest probability, it is also possible that the running time may be different if using a more complicated decoding strategy. However, using a more complex decoding strategy does not affect the results. If a sequence of tokens is memorized, the memorized token will always be the most probable token when generating them.
B.1 Case Study
We also provide a case study of the model’s prediction on the test set regarding the prediction of memorization in Figure 8. From this figure, we can see that:
  1. 1.
    Confirming previous experiments, the memorized token is mostly continuous, showing the memorization happens in chunks of sequences rather than individual sequences.
  2. 2.
    In the first example, the model outputs a fully correct prediction that aligns with the actual label, showing the possibility of predicting memorization by utilizing embedding information.
  3. 3.
    In the second example, we see that the model’s prediction does not align with the actual continuation. The model predicts an unmemorization label for the memorized label.
  4. 4.
    In the third example, we can see this unmemorized sequence. The model fully predicted those labels. We also noticed that the probability for the corresponding token is very high, showing that the model is very confident about the prediction.
Figure 9: 410m Model, Step 2
Figure 10: 410m Model, Step 9
Figure 11: 410m Model, Step 16
Figure 12: 2.8b Model, Step 2
Figure 13: 2.8b Model, Step 9
Figure 14: 2.8b Model, Step 16
Figure 15: 12b Model, Step 2
Figure 16: 12b Model, Step 9
Figure 17: 12b Model, Step 16
B.2 Supplementary Data for Embedding Dynamics
We have presented the PCA visualized embedding dynamics in Figure 5. In this section, we provide the actual numbers of both Cosine Similarity and Euclidean distance to further illustrate this point.
Similarly to Figure 5, we also provide detailed cosine similarity and distance results from Figure 9 to Figure 17. From those figures, we can see that the cosine similarity fluctuates but remains relatively stable for sentences with different memorization scores. Additionally, sentences that are not exactly the same but close regarding the memorization score are very close in the embedding space. For example, the fully memorized sentence is also close to sentences with high memorization scores with embedding similarity over 0.9. This shows the possibility that the model is generating paraphrased memorized sequences. However, with the increase in the model size, the cosine similarity decreases. For the Euclidean distance, we can see that the embedding distances have a decreasing trend with the increasing decoding steps, while the mutual Euclidean increases with the model size.
Figure 18: Index Distribution of Memorized Sequences in 12b model.
B.3 Does LLM prefer to memorize specific parts within the training data?
In this section, we discuss the question of whether LLM prefers to memorize a specific part within the training data. We split the corpora into 50 parts based on their index and examine how many memorized sentences are in those parts.
From the result shown in Figure 18, though the number of memorized sentences is not completely evenly distributed, we can see that there is no significant part that the number of memorized sentences is clearly more than others. This shows the training order does not affect memorization.