### v4

# A Multi-Perspective Analysis of Memorization in Large Language Models

Abstract
Memorization, meaning that Large Language Models (LLMs) can generate content used to train them, is one of the special behaviors of LLMs. Through studies by previous research, little attention has been paid to what makes sentences memorized, how model size affects it, and the dynamics of generating it. In this study, we discuss memorization from multiple perspectives, including scaling model size, input and output dynamics, and less unmemorized content, and reveal: (1) The inter-correlation between memorized/unmemorized sentences, model size, continuation size, and context size, as well as the transition dynamics between sentences of different memorization scores. (2) A boundary effect when generating memorized/unmemorized content and its relation to model size. (3) Sentences of different memorization scores cluster in the embedding space are observed through an embedding dynamics analysis. Entropy analysis in decoding also shows a boundary effect when generating memorized/unmemorized content, albeit oppositely. (4) The possibility of predicting memorization and its relation to model size and continuation length. Further analysis showed that unmemorized sequences are easier to predict than memorized content, which can be explained by the significance of the boundary effect.
1 Introduction
Large language models (LLMs), spanning from BERT Devlin et al. (2019) to the GPT-4 OpenAI et al. (2024), have revolutionized the Natural Language Processing (NLP) and the artificial intelligence research field. Though surprised by their performance on various tasks, we still know little about the underlying mechanism of LLMs whose structures are based on neural models known as black-box Alain and Bengio (2018). Additionally, due to its unprecedented model size and pre-train data size, LLMs start to present unique behaviors Wei et al. (2022), and one of them is memorization.
Memorization Hartmann et al. (2023) in the context of LLMs means the LLM can generate the same content recorded in their pre-train corpus under certain contexts. On the one hand, utilizing memorization, we can use LLMs as a knowledge base Petroni et al. (2019) that can be served for multiple uses. On the other hand, personal information contained in the pre-train corpora may also be elicited maliciously Yao et al. (2024). Previous research Tirumala et al. (2022); Carlini et al. (2023); Biderman et al. (2023b) have studied memorization at the macro level, leaving more micro yet important questions left under-explored, e.g., what makes sentences memorized, what role does the model size play, the input and output dynamics while generating memorized or unmemorized content and how it is related to less memorized content. To answer the above questions, we conducted a multi-perspective study relating memorization to factors like the scaling model size, the input and output dynamics, general statistics, and the prediction of memorization, and we found that:
  1. (I)
    For both memorized and unmemorized sentences, the increase or decrease in model size follows a non-linear trend, indicating a maximum capacity for memorization. The number of memorized sentences decreases sub-linearly with continuation size and increases super-linearly with context size. Moreover, transitions between sentences of different memorization scores were observed, showing that only limited sentences will be memorized with the increase in model size while most sentences remain unmemorized.
  2. (II)
    We conducted an input dynamics study focusing on the n-gram frequency of inputs. A boundary effect was observed when the model began generating memorized/unmemorized content: the frequency of generating unmemorized tokens suddenly decreased while increased for memorized tokens. This boundary effect is also observed at the sentence level and is more obvious in smaller models, indicating that the boundary effect’s significance determines the ease of memorization for a sentence.
  3. (III)
    Additionally, we analyzed decoding dynamics by examining entropy over vocabulary and the drift of decoded embeddings. Entropy analysis revealed an inverse boundary effect, where entropy suddenly increases for unmemorized content and decreases for memorized content. The embedding dynamics analysis showed sentences with different memorization scores cluster in the embedding space, where the mutual embedding distance grows with model size The close distance between sentences with close memorization scores also suggests the existence of paraphrased memorization.
  4. (IV)
    Finally, we trained a Transformer model to predict memorization based on context, showing the possibility of predicting memorization. Additionally, the results suggest that predicting memorization is easier in large models and easier when predicting unmemorized content, which can be explained by the significance of the boundary effect.
2 Related Works
2.1 Scaling Laws of LLM
In this study, our experiments span across various model sizes, thus relating to the research of Scaling Laws Kaplan et al. (2020); Abnar et al. (2021); Villalobos (2023), which suggests the LLM’s performance scales with the size of corpora, the parameter size and the computation required.
Those researches inspired LLMs to continuously scale those factors to gain higher performance like T5 Raffel et al. (2023), GPT-3 Brown et al. (2020), PaLM 2 Anil et al. (2023), etc. On the other hand, researchers also analyzed how the scaling affects particular tasks rather than general performance. For example, Sun and Miceli-Barone (2024) discussed how scaling affects translation tasks and prompt injection attacks. Within those discussions, the LLM’s emergent abilities Wei et al. (2022) are also discovered, which means the LLM suddenly reaches high performance on previous low-performance tasks when the LLM reaches a specific model size. Recent studies have questioned whether emergent abilities are just mirages caused by misuse of metrics Schaeffer et al. (2023) or whether emergent abilities are just context-learning Lu et al. (2023).
Regarding scaling in the field of memorization, Carlini et al. (2023) discussed the memorization across model size and found that the number of memorized texts grows with the model size and context size. Additionally, Biderman et al. (2023a) also discussed similar topics and found that a large portion of memorized text in a small-size model is also memorized by a larger model, showing that memorized texts may share certain common features.
2.2 Memorization
Before LLM and in real-world practices, over-fitting is close to memorization Tetko et al. (1995), which means a near-zero loss in the train set that suggests the neural model perfectly memorized the relation between input and data label Zhang et al. (2017). However, memorization is different from over-fitting as LLM has good performance in the real world while overfitting is usually accompanied by low performance in the test set. Feldman (2020) has conducted a theoretical analysis in classification models explaining why memorization is actually needed. They showed that in a long-tail distribution where many categories only have a few samples, the neural model cannot extract general features from them. The best choice for the neural model is just to memorize them and compare them with data in the test set.
Figure 1: Memorization and Research Scope in this study
For LLMs, different from classification models, they directly generate their pre-train content making the memorization directly observable. which in one way can be used to form a knowledge graph Petroni et al. (2019); AlKhamissi et al. (2022) but in another way also leads to the concern of data contamination Sainz et al. (2023) and privacy risk Yao et al. (2024). Previous research has discussed memorization from a macro level, Tirumala et al. (2022) showing large models tend to memorize samples more easily in the training phase. Carlini et al. (2023) discussed memorization behavior in LLM regarding their factors of model size, continuation size, and context size, showing their inter-correlations. Biderman et al. (2023a) studied memorization in LLM regarding their training phase and the overlap between different model sizes, showing the existence of commonly memorized sentences.
Compared to the previous memorization studies, our research focuses on more micro and under-explored questions, e.g., the dynamics while generating them, the reason why some sentences are memorized, and the prediction of memorization.
3 Experiment Seeting
3.1 Criteria
3.1.1 Memorization Criteria
We choose the K-extractability Carlini et al. (2021) to define the memorization and calculate the memorization score based on it.
We prompt the LLM with a sequence of context tokens $C=\{c_{0}\cdots c_{n}\}$ whose length is defined as context length and use greedy decoding to generate the continuation tokens based on the prompted context. Then, we compare the predicted continuations with the actual continuations and calculate a memorization score using the following equation:
$M(X,Y)=\dfrac{\sum_{i=0}^{n}x_{i}==y_{i}}{L(Y)}$ (1)
[…]
3.1.2 Prediction Criteria
This study uses two criteria to evaluate model performance in predicting memorization, e.g., Token-Level Accuracy and Full Accuracy.
The prediction for a sequence of generated tokens is $\hat{X}=\{\hat{x_{0}},\hat{x_{1}}\dots\hat{x_{n}}\}$ where each prediction $\hat{x_{n}}$ is a binary label indicating the token at this index is a memorized token or not. Similarly, the gold result is denoted as $\hat{Y}=\{\hat{y_{0}},\hat{y_{0}}\dots\hat{y_{n}}\}$ where each $\hat{y}$ is the golden label.
The Token Level Accuracy can be defined as:
$T(\hat{X},\hat{Y})=\dfrac{\sum_{i=0}^{n}\hat{x_{i}}==\hat{y_{i}}}{L(\hat{Y})}$ (2)
The Full Accuracy means that the prediction of token level at a sentence must be entirely correct, which means $T(\hat{X},\hat{Y})==1$.
3.2 Model Setting
We use the Pythia model Biderman et al. (2023b) to analyze the memorization behavior as Pythia provides LLMs trained across various sizes with the same training order using open-sourced Pile Gao et al. (2020) corpora, which benefits the experiment and provides stability.
Specifically, we investigate the model size of [70m, 160m, 410m, 1b, 2.8b, 6.9b, 12b] where $m$ stands for million and $b$ stands for billion. We choose the model trained on the deduplicated pre-train data version to avoid the effect of duplicated sentences, as previous research reports that the chance to be memorized grows exponentially with the number of duplicates Kandpal et al. (2022). [^2] For more details, please refer to https://github.com/EleutherAI/pythia
Figure 2: Memorization Statistics Across Model Size, Complement Size, and Contenxt Size
4 Experiment Results
4.1 Memorization Factors
This section discusses how memorization is affected by different factors. We collect the number of sentences memorized across different extend under different model sizes. Then, we divide them into ten sets with a memorization difference of 0.1. The results are plotted in Figure 2.
Firstly, we can see that the memorized sentences increase with the model size and context size but decrease with the increase of complement size, which aligns with previous research Carlini et al. (2023). We illustrate a more in-depth analysis of those factors in the following sections.
4.1.1 The Factor of Model Size
In this experiment, we discuss how model size affects the number of memorized and unmemorized contexts. From the Figure 2 (a), we can obtain that:
  1. (I)
    The number of sentences with low memorization scores (0-0.3) is significantly higher than those with high memorization scores, indicating that most of the pre-train data are not memorized despite the existence of memorization in LLMs.
  2. (II)
    Among sentences with high memorization scores, the count of fully memorized sentences increases more rapidly, suggesting that LLMs have a propensity to fully memorize sentences rather than partially. Additionally, the number of sentences with low memorization scores (0, 0.1) decreases as the model size increases, indicating that unmemorized sentences gradually become memorized with larger models.
  3. (III)
    The increase or decrease in the number of memorized or unmemorized sentences is not linear with respect to model size. There is a noticeable increase in numbers for fully memorized sentences from 70 million to 2.8 billion parameters, compared to 2.8 billion to 12 billion parameters. A similar decreasing trend for unmemorized sentences is observed. This suggests a capacity for memorization, implying LLMs cannot memorize the entire corpus even with sufficiently large model sizes.
4.1.2 Context and Complement Size
We change the number of prompted tokens while fixing the complement or context size and then observe the number of fully memorized sequences. From the Figure 2 (b) and (c), we can tell that:
  1. (I)
    The decrease in the number of memorized sentences with increasing complement size is not linear. For instance, the complement size increase from 64 to 96 results in a relatively minor decrease compared to the change from 32 to 48, indicating some sentences are firmly memorized.
  2. (II)
    The reduction in memorized sentences with increased complement size is more obvious in larger models. This demonstrates that although larger models memorize more sentences, their memorization is less robust compared to smaller models.
  3. (III)
    The increase in memorized sentences with context size is also non-linear, with longer context leading to an almost exponential rise in the number of memorized sentences. This increase is more significant in larger models, indicating that more contents are potentially memorized in large models, which can be elicited by giving longer context.
Figure 3: Transition Across Different Model Size
4.2 Memorization Transition
This section discusses how sentences with different memorization scores transit across model sizes. Specifically, when trained with a larger model, we study the mutual transition between sentences with different memorization scores. We classify sentences with memorization scores with a range difference of 0.2 with labels of very low, low, medium, high, and very high. We plot the transition matrix in Figure 3, which shows the transition from the 410m size model to the 2.8b size and 2.8b size to the 12b size model, and we can conclude:
  1. (I)
    Most sentences remain in their previous state even when trained with a larger model, as indicated by the diagonal entries in the transition matrices. Additionally, the higher the memorization score, the less likely the sentence will transition to another state. For highly memorized sentences, over 90% remain memorized compared to those with low memorization scores.
  2. (II)
    With increasing model size, sentences are more likely to stay in their original state. For example, the transition probability at the diagonal is higher when moving from the 2.8 billion to the 12 billion parameter model compared to the transition from the 410 million to the 2.8 billion parameter model. This suggests memorized or unmemorized states become more fixed as the model size increases.
  3. (III)
    Even for highly memorized sentences, there is a small chance of transitioning to a low memorized state, implying that some sentences may be memorized randomly rather than due to specific features. This randomness allows for the possibility of transitioning to a low memorized state when trained on a larger model.
4.3 Input Dyanamics
In this section, we discuss the question of whether there is any sign when the model starts to generate memorized or unmemorized content. Especially what makes sentences memorized at different extents and why some sentences are memorized by large-size models but not small-size models.
4.3.1 Token Level Frequency Analysis
Figure 4: One-gram Analysis at Each Index
Firstly, we begin with the input level by analyzing the n-gram statistics of the pre-train corpora. We show the 1-gram statistics across steps in Figure 4, and we can see:
  1. (I)
    A clear boundary effect is observed. Around index 32, representing the first generated token, the frequency drops and then rises for memorized sentences (positive boundary effect), whereas it rises and then drops for unmemorized sentences (negative boundary effect). The negative boundary effect is more pronounced than the positive one.
  2. (II)
    For the half-memorized sentence, we see the frequency increase and drop ((negative boundary effect)) around the index of 39, which is half the length of generated tokens. This shows that for the half-memorized sentence, the previous half is mostly memorized, and the later half is mostly unmemorized, meaning that the memorized tokens are distributed in a near continuous way rather than scattered in the generated tokens.
  3. (III)
    The positive boundary effect in memorized sentences suggests that memorization is driven by the higher frequency of initial tokens, implying that remembering the first few tokens makes the entire sentence easier to memorize. Conversely, the negative boundary effect in unmemorized sentences indicates that the low frequency of initial tokens makes the following sequences easier to forget.
4.3.2 Sentence Level Frequency Analysis
Given the existence of the boundary effect at the token level, we discuss how such an effect exists across model sizes and whether other factors influence memorization at the sequence level. We calculate the average frequency of context and continuation tokens and frequency difference at the boundary in Table 1, and we can obtain:
  1. (I)
    The frequency of one-grams is significantly higher than that of two-grams, approximately 3.5 times higher on average in both context and continuation. The boundary effect is consistent in the two-gram setting, though the positive boundary effect is less obvious due to the frequency drop when computed with two-grams. Despite this, the actual frequency gap remains substantial (million level) when considering the unit is billion.
  2. (II)
    In memorized sentences, the frequency is lower in the context and higher in the continuation, whereas unmemorized sentences exhibit the opposite pattern. This suggests that the boundary effect persists at a broader sequence level, although it is less obvious.
  3. (III)
    For half-memorized sentences, the frequency of continuation tokens is higher than the context average frequency. This is due to the frequency increase before reaching the first generated unmemorized tokens, as indicated in Figure 4.
  4. (IV)
    In the Boundary Frequency Difference column, we see a decreasing trend in both positive and negative boundary effects. However, the decrease in the positive boundary effect makes it less significant, while the decrease in the negative boundary effect makes it more significant. These results imply that the significance of the positive boundary effect correlates with the ease of memorizing a sentence. In contrast, the significance of the negative boundary effect correlates with the ease of not memorizing a sentence.
Size Context Continuation Bounrady Freq Difference
One Two One(B) Two One Two
M H U M H U M H U M H U M H U M H U
160m 1.708 1.713 1.744 0.551 0.534 0.691 1.739 1.837 1.628 0.535 0.659 0.567 0.114 0.330 -0.939 0.033 0.101 -0.663
1b 1.713 1.711 1.752 0.558 0.552 0.697 1.736 1.832 1.631 0.509 0.682 0.564 0.103 0.270 -0.981 0.028 0.090 -0.696
6.8b 1.721 1.710 1.759 0.570 0.565 0.701 1.736 1.829 1.638 0.496 0.699 0.564 0.090 0.140 -0.963 0.027 0.085 -0.726
12b 1.721 1.720 1.760 0.572 0.569 0.702 1.736 1.846 1.626 0.493 0.704 0.563 0.039 0.237 -1.016 0.026 0.083 -0.732
Table 1: Gram Statistics. The frequency unit is billion. Boundary Freq Difference means we use the frequency of the generated token to subtract the frequency of the last token in the context.
4.4 Output Dynamics
4.4.1 Embedding Dynamics
In this section, we discuss the embedding of generated tokens for sentences with different memorization scores. We collect the hidden state at the last layer of each generated token for sentences with different memorization scores and compute the pair-wise Euclidean distance and cosine similarity between them. Then, we perform PCA F.R.S. (1901) on those embeddings and visualize as shown in Figure 5, and we can obtain:
Figure 5: Embedding Dynamics of Across Different Model Size. Memorized Token Num $x$ means $x$ generated tokens are the same with the true continuation.
 [^3] Detailed numbers of Cosine Similarity and Euclidean distance between sentences with different memorization scores are at different decoding steps are at Appendix B.2
  1. (I)
    The cosine similarity remains relatively stable across different encoding steps between sentences memorized at different extents. Meanwhile, from the movement of those points, we can see the Euclidean distance decreases with the generation of tokens. Combined with the previous observation that sentences memorized with different extend have a stable cosine similarity, we can conclude that with the generation of tokens, the angle between sentence vectors at high dimension space remains stable while the cosine similarity decreases.
  2. (II)
    For highly memorized sentences, even when not fully memorized, they cluster closely in the embedding space. This suggests the generated content is both semantically and lexically similar to the memorized content, indicating the existence of paraphrased memorization.
  3. (III)
    Larger models exhibit higher Euclidean distances and lower cosine similarities. The higher Euclidean distance is due to the expansion of hidden sizes (e.g., 512 for 70 million, 2048 for 1 billion parameters). The decrease in cosine similarity is also attributed to the expanded hidden sizes, which increase the expressivity of the embedding space. This expansion enlarges the sentence embeddings, reducing cosine similarity but enhancing the differentiation between sentences. This expanded expressivity in embedding helps to explain the performance gap between different model sizes: larger models distribute different sentences more distinctly with fewer embedding overlaps, while smaller models mix embeddings more, leading to ambiguity and degraded performance.
4.4.2 Generation Dynamics and Entropy
Figure 6: Averaged Entropy at Each Index
This section discusses how the model generates the memorized sentences with different memorization scores, e.g., the generation dynamics. Given the observed boundary effect at the input level, we aim to address whether the model exhibits specific behaviors when generating memorized versus unmemorized sentences. To investigate this, we collected the generation probability for each token during the generation of memorized, unmemorized, and half-memorized sentences. We then calculated the entropy across 10,000 sentences and plotted the entropy for each token in Figure 6. From this analysis, we derive the following observations:
  1. (I)
    First, we can see that the entropy differs depending on the extent of memorization and model size. An unmemorized sentence has a higher average entropy at each token than an unmemorized sentence. This shows LLM is more confident when generating memorized content.
  2. (II)
    Additionally, when the model starts to generate unmemorized tokens, we can see a significant increase in entropy, showing such boundary effect also exists in the decoding steps while behaving oppositely compared to the boundary effect in frequency analysis. The entropy drops around the index of 32 for the memorized and half-memorized content. This also shows the existence of the boundary effect for memorized content. However, we can see the drop of entropy for the memorized sentence is not samely significant as the half-memorized sentence, showing that even for the memorized content, the model is more confident when the sentence is fully memorized, leading to lower entropy.
  3. (III)
    We can also see the entropy decrease with the increase in model size. This suggests that larger models are more confident about generating tokens than small ones. Additionally, we can see that the entropy is lower in the context of memorized tokens, showing that the entropy of context also relates to how a sentence is memorized. Further, the significance of the boundary effect in entropy decreases with the model size for memorized content while remaining unchanged for unmemorized content.
Length 70m 410m 1b 2.8b 6.9b 12b
Token Full Token Full Token Full Token Full Token Full Acc Token
16 78.2 10.2 78.6 10.4 78.8 10.6 80.1 10.7 77.4 8.3 80.3 10.9
32 78.6 5.9 79.6 6.0 79.7 6.1 80.1 6.3 80.5 6.4 80.8 6.4
48 79.6 5.2 80.3 5.4 80.4 5.6 80.4 5.5 80.8 5.8 81.0 6.0
64 80.1 4.7 80.8 4.8 81.2 5.2 81.5 5.5 81.8 5.8 82.1 6.0
Table 2: Performance of Transformer Model on Prediction of Memorization. The context length is defaulted at 32. Length means the token length that is required to be predicted. Token and Full represent Token-Level Accuracy and Full Accuracy. The best results across model sizes are bold, and the best results within a model size are underlined.
4.5 Prediction of Memorization
In this section, we train a Transformer model to predict the memorization of tokens from the LLM’s embedding and statistic features and analyze the results. We discuss whether Transformer can do this task and the difficulty of predicting the memorized token at different model sizes. As in the real world and indicated by our experiment results in previous sections, similar context sentences can also trigger memorized texts Stoehr et al. (2024), and the generated token can also be a paraphrase of the actual continuation Ippolito et al. (2023).
4.5.1 Results on Memorization of Prediction
The training is formalized by receiving both the last layer’s embeddings at each step and statistics (e.g., the entropy). Then, the Transformer is required to generate a binary classification label at each index, predicting whether the token at this index is memorized. The results are presented in Table 2. From the table, we can obtain the following results:
  1. (I)
    First, regarding Token-level accuracy, we can see with a naive Transformer model, the token-level accuracy can reach $80\%$ and even higher, showing that the prediction of memorization at the token level is easy. The full accuracy is low as this requires a correct prediction for every token.
  2. (II)
    As the model size of LLMs increases, both token-level and full-level accuracy improve. This suggests that predicting memorization behavior is easier for larger models because the greater embedding distances make classification easier.
  3. (III)
    Token-level accuracy increases with the continuation size, likely due to increased training data. For instance, a prediction length of 64 involves four times more training data than a prediction length of 16. However, full-sentence accuracy decreases as the prediction length increases, making it harder to achieve full accuracy because more tokens need to be predicted correctly.
4.5.2 Analysis of Full Accuracy
Figure 7: Distribution Across Model Szize of Full Accurate Predictions
In this experiment, we analyze how fully correct predictions distribute across memorization scores to discuss whether the difficulty in predicting memorization changes with the memorization score. From the results shown in Figure 7, we can tell:
  1. (I)
    For any model size, they are more good at predicting sentences with low memorization scores. In contrast, the low portion for sentences with high memorization scores indicate they are harder to predict accurately.
  2. (II)
    As model size increases, the proportion of low memorization scores raise and decrease for high memorization score sentences which even reaches zero for 6.9b model. This suggests larger models are more accurate in predicting unmemorized sentences compared to memorized ones.
  3. (III)
    The ease of predicting sentences with low memorization scores can be attributed to the significance of the boundary effect. Previous experiments show that the boundary effect of unmemorized sequences is more obvious than that of memorized sequences in both token frequency and entropy. Particularly, the boundary effect decreases for memorized sequences while increases for unmemorized sequences with increased model size. Such change makes the boundary effect less significant for memorized sentences, making them harder to predict, while more significant for unmemorized sentences, making them easier to predict.
4.6 Conclusion
In this study, we comprehensively discussed LLM memorization from various perspectives. From a statistical level, our analysis extended the scope of previous research to sentences with lower memorization scores. This is followed by an experiment showing the memorization transition across model sizes. At the input dynamic level, through frequency analysis, we found the positive and negative boundary effects when generating memorized and unmemorized tokens and their relation to how easily a sentence can be memorized/unmemorized. In the output dynamics, at the embedding level, we found clusters of sentences with different memorization scores in the embedding space, and the close distance of sentences with high memorization scores indicates the existence of paraphrase memorization. At the entropy level, we observed an opposite boundary effect and analyzed its change with model size. Finally, we trained a Transformer model to predict the memorization, showing that token-level prediction is easy while sentence-level is challenging. Through analysis of fully correct predicted samples, we found unmemorized tokens are easier to predict than memorized tokens, which can be explained by the significance of the boundary effect.
5 Limitations
This research has analyzed various factors regarding the memorization phenomena of LLMs. We acknowledge that there are still limitations to this research. Due to the lack of LLMs whose models and data are both being released, as the data is also different, it is hard to compare the memorization across different LLMs. Future research can be on how to comprehensively measure memorization of various LLMs, either close-sourced like GPT-4 or open-sourced like Pythia or LLaMa. Additionally, Pythia only provides LLMs up to 12B, which is still smaller than SotA open-sourced LLMs like LLaMA, whose largest model is 70B. The difference between 12B and 70B may lead to emergent abilities, which may also affect the memorization results obtained in 12B, which may not be able to scale even further due to those factors. Though we trained the Transformer to predict the memorization of LLMs, it is more analysis-oriented; thus, the performance is not the main focus of this experiment. Additionally, in this research, the discussion of memorization is under verbatim memorization, where generated tokens are identical to the same sentence in the corpora.
Appendix A Appendix
Figure 8: Prediction Examples. Pred Prob means the output predicted probability of the corresponding label in the pred row for each example. Gold means the true label, Pred means the predicted label, and Pred Prob means the probability of the corresponding prediction. $M$ means the label for a token at this index is a memorized token. $U$ means the label for a token at this index is an unmemorized token.
Appendix B Experiment Setting
The experiment uses 64 A100 40Gbs GPUs when using LLMs to generate tokens given the previous context, which utilizes PyTorch’s parallel running packages. We run the model with half-precision, which increases both the speed and saves memory. This follows the previous Pythia implementation when generating tokens given context.
The running time depends on the model size and the generated token length. In the situation when a 70m model with 32 context tokens and 16 tokens is required to be generated, it can be run with one A100 GPU within several hours. However, if such an experiment is in a single A100 GPU, it would estimated to take two weeks to finish the generation. Therefore, with 64 GPUs, the running of the 12b model in such a situation can be shortened to around one day. However, the generation time also largely increases with the length of generated tokens, which grows in a linear way with the number of tokens to be generated. Additionally, since we use greedy decoding and do not consider other possible decoding options but the token with the highest probability, it is also possible that the running time may be different if using a more complicated decoding strategy. However, using a more complex decoding strategy does not affect the results. If a sequence of tokens is memorized, the memorized token will always be the most probable token when generating them.
B.1 Case Study
We also provide a case study of the model’s prediction on the test set regarding the prediction of memorization in Figure 8. From this figure, we can see that:
  1. 1.
    Confirming previous experiments, the memorized token is mostly continuous, showing the memorization happens in chunks of sequences rather than individual sequences.
  2. 2.
    In the first example, the model outputs a fully correct prediction that aligns with the actual label, showing the possibility of predicting memorization by utilizing embedding information.
  3. 3.
    In the second example, we see that the model’s prediction does not align with the actual continuation. The model predicts an unmemorization label for the memorized label.
  4. 4.
    In the third example, we can see this unmemorized sequence. The model fully predicted those labels. We also noticed that the probability for the corresponding token is very high, showing that the model is very confident about the prediction.
Figure 9: 410m Model, Step 2
Figure 10: 410m Model, Step 9
Figure 11: 410m Model, Step 16
Figure 12: 2.8b Model, Step 2
Figure 13: 2.8b Model, Step 9
Figure 14: 2.8b Model, Step 16
Figure 15: 12b Model, Step 2
Figure 16: 12b Model, Step 9
Figure 17: 12b Model, Step 16
B.2 Supplementary Data for Embedding Dynamics
We have presented the PCA visualized embedding dynamics in Figure 5. In this section, we provide the actual numbers of both Cosine Similarity and Euclidean distance to further illustrate this point.
Similarly to Figure 5, we also provide detailed cosine similarity and distance results from Figure 9 to Figure 17. From those figures, we can see that the cosine similarity fluctuates but remains relatively stable for sentences with different memorization scores. Additionally, sentences that are not exactly the same but close regarding the memorization score are very close in the embedding space. For example, fully memorized sentences are also close to sentences with high memorization scores and embedding similarity of over 0.9. This shows the possibility that the model is generating paraphrased memorized sequences. However, with the increase in the model size, the cosine similarity decreases. For the Euclidean distance, we can see that the embedding distances have a decreasing trend with the increasing decoding steps, while the mutual Euclidean increases with the model size.
Figure 18: Index Distribution of Memorized Sequences in 12b model.
B.3 Does LLM prefer to memorize specific parts within the training data?
In this section, we discuss the question of whether LLM prefers to memorize a specific part within the training data. We split the corpora into 50 parts based on their index and examine how many memorized sentences are in those parts.
From the result shown in Figure 18, though the number of memorized sentences is not completely evenly distributed, we can see that there is no significant part that the number of memorized sentences is clearly more than others. This shows the training order does not affect memorization.
B.4 Detail Gram Statistics
We also provided detailed 1-gram, 2-gran, and 3-gram statistics for the memorized, un-memorized, half-memorized, and quarter-memorized contents. The average frequency is calculated at each step, and we compute the average frequency of the context, the average frequency of the continuation, and the average frequency of the whole sequence. The results span from 70m size model to 12b size model shown in Table 3 to Table 6.
target memorized memorized memorized memorized memorized memorized memorized forgotten forgotten forgotten forgotten forgotten forgotten forgotten
size 70m 160m 410m 1b 2.8b 6.9b 12b 70m 160m 410m 1b 2.8b 6.9b 12b
0 1,761,877,325 1,744,677,836 1,753,317,716 1,749,565,941 1,759,384,624 1,763,400,000 1,794,900,000 1,732,531,426 1,732,526,814 1,733,883,954 1,735,430,248 1,757,800,000 1,717,300,000 1,753,200,000
1 1,756,407,896 1,733,466,452 1,737,641,504 1,739,836,645 1,747,576,577 1,712,000,000 1,742,500,000 1,733,790,778 1,734,398,485 1,735,811,740 1,737,110,533 1,754,500,000 1,724,000,000 1,759,900,000
2 1,754,544,442 1,735,828,900 1,739,272,135 1,740,794,079 1,750,863,001 1,758,000,000 1,804,000,000 1,733,955,892 1,734,433,227 1,735,927,786 1,737,263,808 1,719,800,000 1,730,300,000 1,755,600,000
3 1,745,749,714 1,740,145,636 1,739,384,453 1,740,237,878 1,749,816,871 1,700,300,000 1,818,500,000 1,734,815,677 1,735,031,658 1,736,843,088 1,737,987,235 1,686,300,000 1,765,900,000 1,779,800,000
4 1,733,568,727 1,729,484,889 1,728,938,412 1,733,592,149 1,742,113,248 1,766,400,000 1,708,800,000 1,735,128,356 1,735,452,086 1,737,120,799 1,738,793,702 1,767,300,000 1,733,300,000 1,759,200,000
5 1,727,964,181 1,723,277,679 1,725,427,437 1,727,321,616 1,736,693,521 1,738,500,000 1,710,400,000 1,736,270,145 1,736,331,860 1,737,850,119 1,740,340,490 1,758,200,000 1,717,900,000 1,739,900,000
6 1,725,841,351 1,716,563,761 1,723,081,626 1,725,702,469 1,735,415,386 1,758,000,000 1,742,900,000 1,735,559,708 1,736,435,200 1,738,820,970 1,740,165,047 1,780,700,000 1,742,500,000 1,713,400,000
7 1,722,343,456 1,711,523,969 1,719,492,821 1,724,003,560 1,732,499,267 1,698,400,000 1,731,400,000 1,736,139,862 1,737,411,285 1,738,841,087 1,740,979,176 1,708,700,000 1,726,500,000 1,761,400,000
8 1,715,952,971 1,703,541,696 1,712,534,695 1,717,230,465 1,730,417,914 1,679,100,000 1,709,000,000 1,736,938,342 1,738,000,206 1,739,984,548 1,741,715,014 1,752,400,000 1,775,600,000 1,740,500,000
9 1,701,602,676 1,696,501,612 1,705,597,804 1,710,036,650 1,721,980,996 1,776,400,000 1,716,900,000 1,737,218,230 1,738,548,131 1,740,764,383 1,742,163,420 1,759,400,000 1,722,800,000 1,764,800,000
10 1,708,709,485 1,703,507,469 1,707,974,902 1,712,561,959 1,724,060,121 1,672,300,000 1,733,400,000 1,737,059,426 1,738,812,313 1,740,973,658 1,742,911,460 1,760,200,000 1,807,900,000 1,753,800,000
11 1,708,002,048 1,700,334,479 1,705,745,789 1,712,145,483 1,721,887,809 1,719,300,000 1,729,500,000 1,737,667,494 1,739,075,343 1,741,033,716 1,742,470,266 1,743,600,000 1,812,600,000 1,758,700,000
12 1,704,765,267 1,690,528,200 1,700,680,032 1,703,727,332 1,714,582,320 1,739,700,000 1,711,800,000 1,737,833,800 1,738,990,635 1,741,004,691 1,742,710,814 1,768,300,000 1,754,700,000 1,771,800,000
13 1,693,927,960 1,683,230,446 1,695,709,392 1,702,676,948 1,712,420,697 1,731,000,000 1,690,500,000 1,737,172,555 1,738,180,555 1,740,335,149 1,741,471,936 1,698,100,000 1,745,800,000 1,745,100,000
14 1,696,634,570 1,681,852,317 1,695,295,636 1,701,256,215 1,712,257,191 1,725,700,000 1,699,700,000 1,737,522,088 1,738,958,841 1,740,701,751 1,742,059,832 1,740,500,000 1,699,000,000 1,728,500,000
15 1,679,551,926 1,671,142,177 1,688,416,481 1,697,105,648 1,709,669,411 1,752,700,000 1,720,700,000 1,736,261,928 1,737,987,431 1,740,131,961 1,741,684,474 1,752,300,000 1,705,000,000 1,746,200,000
16 1,667,903,751 1,670,413,255 1,678,747,207 1,691,187,011 1,704,670,219 1,719,900,000 1,699,800,000 1,735,533,670 1,737,210,612 1,738,579,355 1,740,344,603 1,750,300,000 1,781,300,000 1,752,600,000
17 1,682,336,989 1,679,636,457 1,684,708,172 1,696,051,548 1,706,414,493 1,718,300,000 1,727,400,000 1,736,647,826 1,737,903,626 1,739,991,371 1,741,869,751 1,754,200,000 1,752,800,000 1,780,200,000
18 1,696,024,732 1,688,157,613 1,693,550,892 1,701,124,435 1,711,161,137 1,689,300,000 1,685,300,000 1,736,875,723 1,738,169,859 1,740,186,111 1,742,150,532 1,723,700,000 1,764,500,000 1,720,000,000
19 1,700,044,430 1,691,379,251 1,700,787,282 1,703,484,170 1,713,713,338 1,735,600,000 1,698,900,000 1,736,115,268 1,738,516,054 1,740,906,141 1,742,896,568 1,716,700,000 1,778,500,000 1,739,500,000
20 1,710,839,951 1,700,701,910 1,706,890,340 1,710,112,768 1,717,662,092 1,703,300,000 1,750,400,000 1,735,637,929 1,737,519,099 1,740,228,710 1,741,971,117 1,776,900,000 1,761,300,000 1,752,300,000
21 $71$,696,906,569 1,701,929,069 1,706,919,694 1,714,832,769 1,726,500,000 1,722,600,000 1,736,494,732 1,739,238,409 1,742,035,789 1,744,490,773 1,747,300,000 1,730,200,000 1,727,900,000
22 1,707,031,048 1,694,596,312 1,700,606,809 1,707,393,546 1,714,688,122 1,688,200,000 1,719,400,000 1,735,343,187 1,738,714,382 1,741,858,543 1,743,975,469 1,763,800,000 1,738,800,000 1,757,900,000
23 1,720,840,503 1,710,023,803 1,709,653,995 1,713,995,947 1,719,166,026 1,728,600,000 1,726,500,000 1,735,401,486 1,738,931,803 1,743,158,826 1,745,299,877 1,733,700,000 1,745,300,000 1,764,500,000
24 1,731,576,174 1,715,940,520 1,714,003,336 1,715,023,292 1,717,926,947 1,758,400,000 1,722,100,000 1,734,596,940 1,738,386,519 1,742,505,032 1,744,500,841 1,749,300,000 1,761,700,000 1,760,400,000
25 1,752,003,292 1,731,855,788 1,727,114,867 1,724,798,309 1,728,919,793 1,724,100,000 1,708,500,000 1,737,772,973 1,742,462,185 1,746,923,721 1,749,250,474 1,773,200,000 1,764,900,000 1,739,600,000
26 1,755,434,381 1,735,918,727 1,728,805,371 1,726,260,802 1,725,855,889 1,717,700,000 1,710,400,000 1,736,254,689 1,741,221,711 1,746,794,790 1,750,257,624 1,782,600,000 1,737,100,000 1,737,500,000
27 1,751,078,472 1,732,474,044 1,721,053,568 1,718,911,936 1,720,139,844 1,730,200,000 1,697,500,000 1,737,465,109 1,745,018,430 1,752,091,624 1,756,426,168 1,732,400,000 1,741,600,000 1,753,500,000
28 1,752,459,582 1,738,590,437 1,733,534,889 1,725,443,806 1,726,659,853 1,686,300,000 1,679,200,000 1,727,033,252 1,738,018,143 1,747,326,896 1,754,100,491 1,764,800,000 1,752,900,000 1,767,200,000
29 1,727,987,920 1,710,624,126 1,706,925,069 1,698,887,785 1,701,059,926 1,699,000,000 1,699,700,000 1,742,590,726 1,758,725,331 1,772,881,062 1,782,612,036 1,789,600,000 1,790,700,000 1,810,600,000
30 1,723,009,291 1,712,857,933 1,705,443,834 1,699,764,161 1,701,619,618 1,711,900,000 1,710,500,000 1,742,138,296 1,764,672,256 1,781,900,380 1,793,731,882 1,863,100,000 1,825,200,000 1,818,400,000
31 1,697,576,775 1,687,242,321 1,674,099,012 1,668,454,050 1,668,926,625 1,649,200,000 1,683,400,000 1,856,009,443 1,902,117,823 1,941,790,973 1,967,891,756 1,975,400,000 1,987,200,000 2,005,800,000
Avg Context 1,719,510,719 1,708,216,456 1,711,448,892 1,713,925,259 1,721,720,489 1,721,178,125 1,722,078,125 1,739,930,530 1,744,606,260 1,749,349,648 1,752,719,607 1,759,534,375 1,759,221,875 1,763,115,625
32 1,812,724,891 1,801,299,369 1,783,560,981 1,771,127,395 1,765,849,310 1,739,200,000 1,722,000,000 957,760,109 963,964,253 981,205,278 986,644,973 1,017,100,000 1,024,300,000 989,330,000
33 1,785,198,082 1,782,599,042 1,764,708,279 1,754,286,257 1,750,603,678 1,748,500,000 1,734,000,000 1,477,089,688 1,452,367,316 1,435,934,090 1,428,928,547 1,415,800,000 1,418,800,000 1,417,900,000
34 1,772,561,111 1,767,491,083 1,753,649,952 1,744,671,079 1,742,356,129 1,769,300,000 1,735,800,000 1,623,949,166 1,609,532,509 1,597,061,854 1,592,797,469 1,560,600,000 1,576,000,000 1,592,600,000
35 1,762,468,180 1,755,171,008 1,742,357,780 1,734,034,726 1,733,128,170 1,732,800,000 1,735,800,000 1,656,747,957 1,650,264,891 1,644,001,022 1,641,698,771 1,625,200,000 1,654,900,000 1,635,900,000
36 1,751,084,744 1,744,819,021 1,731,757,083 1,727,988,589 1,726,739,343 1,728,600,000 1,706,600,000 1,678,449,622 1,673,912,651 1,669,885,767 1,668,314,472 1,642,100,000 1,683,100,000 1,643,100,000
37 1,737,469,976 1,729,200,256 1,720,955,403 1,720,387,855 1,719,147,209 1,712,900,000 1,787,300,000 1,694,329,815 1,691,870,813 1,689,952,951 1,689,876,998 1,727,300,000 1,699,700,000 1,699,200,000
38 1,737,353,894 1,724,369,474 1,716,962,208 1,710,620,923 1,711,850,544 1,691,900,000 1,761,400,000 1,696,281,632 1,695,900,348 1,694,668,623 1,695,468,826 1,674,600,000 1,697,300,000 1,679,200,000
39 1,715,081,363 1,703,843,951 1,703,301,172 1,705,983,119 1,709,368,578 1,706,400,000 1,660,200,000 1,700,949,641 1,700,918,018 1,700,662,603 1,701,381,229 1,735,200,000 1,703,000,000 1,694,800,000
40 1,686,685,834 1,682,653,925 1,690,744,526 1,694,008,976 1,698,953,947 1,698,400,000 1,732,100,000 1,699,950,424 1,700,320,136 1,700,698,424 1,701,888,972 1,709,600,000 1,749,200,000 1,717,200,000
41 1,697,887,078 1,690,596,776 1,699,039,050 1,702,293,288 1,706,226,135 1,706,000,000 1,729,200,000 1,701,980,579 1,703,673,165 1,704,305,693 1,705,426,679 1,713,800,000 1,688,200,000 1,721,300,000
42 1,706,446,759 1,699,625,915 1,704,257,959 1,710,138,829 1,711,973,193 1,706,800,000 1,696,400,000 1,699,916,580 1,701,963,362 1,704,179,446 1,704,239,602 1,693,500,000 1,680,700,000 1,699,700,000
43 1,724,996,796 1,711,048,851 1,718,405,094 1,721,177,377 1,728,269,651 1,730,100,000 1,737,300,000 1,701,424,868 1,702,896,473 1,704,910,899 1,706,115,083 1,699,300,000 1,718,400,000 1,702,000,000
44 1,729,978,559 1,717,877,194 1,721,748,175 1,727,317,540 1,725,798,564 1,752,400,000 1,680,700,000 1,699,235,317 1,702,231,574 1,704,286,186 1,705,328,377 1,732,700,000 1,719,100,000 1,715,900,000
45 1,756,734,060 1,740,464,386 1,745,501,552 1,750,234,915 1,752,814,925 1,722,000,000 1,740,000,000 1,699,769,885 1,702,350,463 1,703,950,116 1,705,365,109 1,683,000,000 1,753,700,000 1,696,100,000
46 1,772,357,723 1,765,976,491 1,769,251,508 1,775,699,434 1,780,492,152 1,796,400,000 1,773,900,000 1,697,965,233 1,700,236,541 1,702,146,238 1,703,508,158 1,740,600,000 1,727,900,000 1,724,700,000
47 1,826,784,956 1,810,274,933 1,823,557,511 1,829,605,754 1,834,611,860 1,849,800,000 1,853,200,000 1,694,816,806 1,697,715,045 1,699,839,943 1,701,149,553 1,730,400,000 1,716,300,000 1,675,600,000
Avg Continuation 1,748,488,375 1,739,206,980 1,736,859,890 1,736,223,503 1,737,386,462 1,736,968,750 1,736,618,750 1,630,038,582 1,628,132,347 1,627,355,571 1,627,383,301 1,631,300,000 1,638,162,500 1,625,283,125
Avg Sentence 1,729,169,938 1,723,124,945 1,719,919,225 1,721,358,007 1,726,942,480 1,726,441,667 1,726,925,000 1,703,299,881 1,705,781,622 1,708,684,955 1,710,940,838 1,716,789,583 1,718,868,750 1,717,171,458
Table 3: One Gram Statistics for Memorized and Unmemorized Content
target half half half half half half half quarter quarter quarter quarter quarter quarter quarter
size 70m 160m 410m 1b 2.8b 6.9b 12b 70m 160m 410m 1b 2.8b 6.9b 12b
0 1,747,500,000 1,780,200,000 1,752,900,000 1,700,100,000 1,729,500,000 1,733,700,000 1,715,500,000 1774800000 1,754,800,000 1785800000 1763800000 1781400000 1778900000 1756200000
1 1,771,400,000 1,735,900,000 1,707,600,000 1,717,800,000 1,756,600,000 1,731,400,000 1,729,600,000 1746500000 1,781,100,000 1767400000 1746500000 1742600000 1779500000 1736400000
2 1,747,800,000 1,745,400,000 1,758,500,000 1,731,800,000 1,705,000,000 1,687,200,000 1,774,100,000 1754000000 1,794,900,000 1781500000 1756500000 1768900000 1760300000 1753800000
3 1,760,400,000 1,712,100,000 1,743,700,000 1,745,200,000 1,726,600,000 1,753,500,000 1,763,800,000 1791200000 1,776,600,000 1769500000 1776000000 1791900000 1748200000 1782600000
4 1,693,800,000 1,720,700,000 1,726,500,000 1,696,900,000 1,748,200,000 1,697,300,000 1,723,500,000 1696800000 1,740,300,000 1789400000 1761900000 1792900000 1760800000 1754400000
5 1,711,200,000 1,746,400,000 1,762,300,000 1,722,900,000 1,680,700,000 1,715,700,000 1,714,000,000 1761700000 1,765,300,000 1777000000 1720700000 1763800000 1765800000 1746100000
6 1,700,900,000 1,717,100,000 1,673,600,000 1,753,500,000 1,774,500,000 1,735,800,000 1,771,000,000 1764900000 1,757,900,000 1726200000 1744400000 1742400000 1761400000 1785200000
7 1,737,400,000 1,743,900,000 1,710,200,000 1,741,300,000 1,766,200,000 1,688,400,000 1,730,600,000 1755400000 1,690,800,000 1790000000 1717000000 1751600000 1740100000 1754900000
8 1,705,800,000 1,721,300,000 1,728,200,000 1,723,200,000 1,720,900,000 1,676,100,000 1,738,200,000 1739900000 1,726,100,000 1799800000 1773900000 1806700000 1704900000 1737600000
9 1,754,600,000 1,756,100,000 1,762,300,000 1,722,400,000 1,767,400,000 1,772,800,000 1,708,800,000 1785800000 1,741,600,000 1756200000 1763000000 1744900000 1757800000 1708900000
10 1,737,700,000 1,726,600,000 1,761,000,000 1,715,500,000 1,724,600,000 1,754,200,000 1,757,500,000 1753600000 1,763,700,000 1757000000 1760600000 1748400000 1736600000 1770700000
11 1,778,500,000 1,738,600,000 1,691,800,000 1,753,500,000 1,704,100,000 1,744,300,000 1,765,900,000 1774300000 1,773,600,000 1764500000 1746700000 1752300000 1741700000 1740200000
12 1,721,500,000 1,696,300,000 1,678,400,000 1,699,000,000 1,714,100,000 1,718,900,000 1,765,400,000 1776700000 1,734,000,000 1771900000 1716500000 1736100000 1727900000 1730800000
13 1,751,300,000 1,732,400,000 1,708,000,000 1,700,400,000 1,698,800,000 1,754,600,000 1,719,200,000 1714300000 1,744,900,000 1749500000 1741100000 1729200000 1746400000 1742400000
14 1,737,600,000 1,731,900,000 1,735,500,000 1,682,700,000 1,695,200,000 1,701,600,000 1,730,800,000 1751000000 1,749,400,000 1731600000 1773600000 1738200000 1766800000 1712100000
15 1,693,700,000 1,725,700,000 1,774,900,000 1,732,300,000 1,713,300,000 1,775,300,000 1,743,900,000 1728000000 1,724,300,000 1739000000 1742500000 1762500000 1736200000 1782100000
16 1,739,000,000 1,749,900,000 1,740,800,000 1,777,700,000 1,747,300,000 1,749,300,000 1,702,800,000 1738500000 1,740,800,000 1705500000 1771500000 1729500000 1763500000 1749200000
17 1,741,900,000 1,694,200,000 1,742,100,000 1,771,200,000 1,721,200,000 1,741,300,000 1,783,000,000 1745700000 1,740,100,000 1767500000 1736700000 1738900000 1741000000 1761400000
18 1,738,200,000 1,703,300,000 1,740,700,000 1,711,900,000 1,705,700,000 1,704,600,000 1,686,400,000 1737100000 1,743,400,000 1743200000 1737100000 1724500000 1735100000 1756800000
19 1,728,500,000 1,725,600,000 1,729,100,000 1,719,600,000 1,725,700,000 1,679,600,000 1,747,000,000 1726900000 1,791,300,000 1728300000 1740300000 1707300000 1736600000 1719900000
20 1,691,300,000 1,761,400,000 1,714,800,000 1,761,200,000 1,695,900,000 1,702,800,000 1,733,400,000 1727700000 1,767,400,000 1707000000 1730800000 1771400000 1725200000 1795800000
21 1,677,600,000 1,738,600,000 1,699,400,000 1,737,100,000 1,691,400,000 1,680,900,000 1,697,700,000 1755200000 1,741,500,000 1750000000 1777700000 1751000000 1741700000 1746900000
22 1,740,300,000 1,713,600,000 1,716,200,000 1,668,000,000 1,698,100,000 1,718,800,000 1,722,600,000 1774400000 1,710,600,000 1757000000 1774700000 1797700000 1752900000 1746100000
23 1,748,200,000 1,666,100,000 1,664,400,000 1,729,100,000 1,677,100,000 1,734,500,000 1,731,200,000 1747300000 1,737,000,000 1766600000 1782700000 1761900000 1757800000 1723400000
24 1,691,700,000 1,735,300,000 1,716,500,000 1,722,300,000 1,733,100,000 1,699,600,000 1,678,600,000 1735800000 1,734,200,000 1748600000 1712500000 1743000000 1772400000 1700900000
25 1,728,800,000 1,656,600,000 1,726,600,000 1,700,200,000 1,690,500,000 1,697,900,000 1,732,800,000 1731200000 1,766,800,000 1740300000 1724100000 1736800000 1782600000 1718200000
26 1,700,700,000 1,715,200,000 1,654,200,000 1,694,200,000 1,703,700,000 1,684,400,000 1,733,500,000 1708800000 1,715,100,000 1739000000 1734800000 1721100000 1695100000 1749000000
27 1,724,500,000 1,645,400,000 1,754,400,000 1,707,800,000 1,724,900,000 1,721,800,000 1,672,900,000 1703300000 1,715,900,000 1711100000 1697200000 1718800000 1708100000 1695900000
28 1,658,300,000 1,712,500,000 1,744,100,000 1,660,800,000 1,739,500,000 1,678,300,000 1,701,000,000 1743400000 1,732,800,000 1752700000 1712300000 1724300000 1736400000 1680500000
29 1,692,600,000 1,667,500,000 1,686,700,000 1,655,700,000 1,657,100,000 1,676,700,000 1,707,500,000 1697300000 1,771,300,000 1731300000 1717800000 1677500000 1714100000 1691500000
30 1,655,500,000 1,690,700,000 1,658,200,000 1,647,300,000 1,683,200,000 1,639,600,000 1,628,200,000 1760800000 1,689,800,000 1675800000 1691500000 1678700000 1649800000 1697800000
31 1,666,800,000 1,541,100,000 1,637,100,000 1,564,000,000 1,513,400,000 1,579,400,000 1,554,100,000 1562100000 1,515,300,000 1536200000 1484100000 1477500000 1436800000 1507600000
Avg Context 1,721,093,750 1,713,987,500 1,718,771,875 1,711,456,250 1,710,421,875 1,710,321,875 1,720,765,625 1,739,512,500 1,738,518,750 1,744,262,500 1,735,328,125 1,737,928,125 1,733,200,000 1,732,353,125
32 1,878,300,000 1,871,300,000 1,800,300,000 1,834,700,000 1,812,500,000 1,719,500,000 1,791,700,000 2189100000 2,090,900,000 2020500000 1992900000 1907700000 1938700000 1895000000
33 1,842,000,000 1,854,700,000 1,836,200,000 1,829,200,000 1,869,400,000 1,834,500,000 1,749,400,000 2458900000 2,344,000,000 2383700000 2295800000 2224900000 2269700000 2275200000
34 1,842,600,000 1,900,800,000 1,843,500,000 1,862,600,000 1,778,400,000 1,812,100,000 1,832,600,000 2650800000 2,597,500,000 2592700000 2537800000 2483300000 2511000000 2568800000
35 1,841,600,000 1,875,500,000 1,883,700,000 1,894,300,000 1,866,800,000 1,843,400,000 1,868,600,000 2520900000 2,583,300,000 2586100000 2591600000 2586700000 2599000000 2572600000
36 1,871,800,000 1,901,500,000 1,932,600,000 1,931,300,000 1,926,100,000 1,872,300,000 1,893,200,000 1778900000 1,742,400,000 1743700000 1696700000 1687300000 1715400000 1720500000
37 1,885,600,000 1,921,800,000 1,997,900,000 1,915,800,000 1,981,000,000 2,077,100,000 2,038,100,000 1736300000 1,739,900,000 1728100000 1685100000 1664500000 1681500000 1700300000
38 1,998,200,000 2,101,300,000 2,081,700,000 2,115,900,000 2,156,700,000 2,209,100,000 2,254,500,000 1762300000 1,710,100,000 1711100000 1737600000 1718500000 1736000000 1705300000
39 2,032,100,000 2,158,400,000 2,168,100,000 2,189,200,000 2,246,300,000 2,276,200,000 2,369,000,000 1809200000 1,768,100,000 1799700000 1747600000 1744100000 1738300000 1758800000
40 1,723,200,000 1,690,600,000 1,715,300,000 1,692,500,000 1,691,600,000 1,691,900,000 1,719,100,000 1765200000 1,751,200,000 1789500000 1746300000 1757000000 1733900000 1779400000
41 1,738,700,000 1,731,600,000 1,699,700,000 1,709,900,000 1,706,600,000 1,668,700,000 1,706,000,000 1773400000 1,777,100,000 1770200000 1803300000 1780300000 1788700000 1780500000
42 1,707,400,000 1,795,200,000 1,719,700,000 1,721,300,000 1,743,300,000 1,700,900,000 1,725,800,000 1822600000 1,814,600,000 1763300000 1800800000 1793500000 1788700000 1740000000
43 1751700000 1,708,000,000 1,720,500,000 1,713,600,000 1,733,400,000 1,727,600,000 1,687,200,000 1839800000 1,774,600,000 1838300000 1761600000 1754700000 1787300000 1805000000
44 1753600000 1,742,000,000 1,764,800,000 1,694,000,000 1,739,700,000 1,686,500,000 1,721,000,000 1760300000 1,806,500,000 1817700000 1760600000 1787800000 1782700000 1752900000
45 1729900000 1,720,800,000 1,715,200,000 1,721,700,000 1,786,500,000 1,697,200,000 1,733,700,000 1775700000 1,763,900,000 1829800000 1840800000 1784300000 1822300000 1734000000
46 1733000000 1,717,200,000 1,741,500,000 1,705,100,000 1,750,600,000 1,721,000,000 1,737,400,000 1797500000 1,799,100,000 1788600000 1729000000 1772000000 1765100000 1758700000
47 1732500000 1,715,600,000 1,748,500,000 1,788,200,000 1,710,100,000 1,731,500,000 1,712,500,000 1763400000 1,757,300,000 1783200000 1835300000 1791700000 1764800000 1766000000
Avg Continuation 1,816,387,500 1,837,893,750 1,835,575,000 1,832,456,250 1,843,687,500 1,829,343,750 1,846,237,500 1,950,268,750 1,926,281,250 1,934,137,500 1,910,175,000 1,889,893,750 1,901,443,750 1,894,562,500
Avg Sentence 1,752,858,333 1,755,289,583 1,757,706,250 1,751,789,583 1,754,843,750 1,749,995,833 1,762,589,583 1,809,764,583 1,801,106,250 1,807,554,167 1,793,610,417 1,788,583,333 1,789,281,250 1,786,422,917
Table 4: One Gram Statistics for Half-memorized and Quarter-memorized content
target memorized memorized memorized memorized memorized memorized memorized forgotten forgotten forgotten forgotten forgotten forgotten forgotten
size 70m 160m 410m 1b 2.8b 6.9b 12b 70m 160m 410m 1b 2.8b 6.9b 12b
1 56,865,298 55,633,408 57,631,394 57,851,391 58,747,419 59,390,492 59,538,566 67,723,858 67,753,352 67,683,530 67,708,452 67,740,430 67,754,040 67,767,823
2 57,323,714 56,317,639 57,433,612 57,993,010 58,602,181 59,488,016 59,803,389 67,737,009 67,768,452 67,769,364 67,815,489 67,769,957 67,800,825 67,828,380
3 56,730,613 56,231,828 57,388,365 58,009,821 58,928,771 59,525,329 59,598,379 67,775,953 67,738,481 67,807,016 67,780,588 67,804,139 67,851,921 67,894,036
4 55,911,242 56,234,359 56,980,241 57,792,288 58,806,766 59,391,815 59,636,502 67,949,220 67,825,207 67,839,517 67,841,697 67,867,488 67,909,524 67,967,557
5 55,692,756 55,655,983 56,819,836 57,273,741 58,351,556 59,022,366 59,346,451 67,958,033 67,886,128 67,874,628 68,003,000 67,965,687 68,045,776 68,030,203
6 55,894,920 55,666,111 56,911,193 57,438,723 58,686,676 59,320,405 59,557,076 67,985,473 67,869,913 67,917,993 68,012,109 67,974,605 68,003,875 68,007,874
7 56,047,068 55,753,285 56,733,219 57,360,465 58,300,186 59,087,792 59,572,316 68,058,487 68,015,919 68,064,231 68,101,370 68,115,977 68,153,223 68,184,607
8 55,832,397 55,349,442 56,329,311 57,111,974 58,349,971 58,924,645 59,505,541 68,017,225 68,013,215 68,034,561 68,111,358 68,101,007 68,139,029 68,201,839
9 55,253,643 54,713,056 55,885,459 56,532,115 57,837,312 58,610,331 59,086,863 68,135,592 68,114,172 68,143,152 68,214,493 68,138,035 68,189,743 68,155,354
10 54,503,716 54,257,056 55,495,149 56,166,844 57,307,569 58,022,654 58,377,791 68,160,293 68,172,545 68,187,816 68,250,398 68,215,884 68,267,972 68,242,452
11 54,988,987 54,708,110 55,773,288 56,333,064 57,701,866 58,500,535 58,871,367 68,234,320 68,217,765 68,269,026 68,248,385 68,243,497 68,268,692 68,241,732
12 54,935,679 54,664,019 55,533,124 56,267,162 57,079,217 58,003,846 58,331,501 68,213,578 68,234,560 68,252,214 68,247,968 68,328,780 68,343,161 68,299,170
13 54,767,327 54,145,183 55,296,145 55,857,794 56,871,961 57,652,393 57,921,174 68,226,649 68,274,620 68,277,162 68,277,302 68,337,296 68,382,126 68,389,974
14 54,021,320 53,569,285 54,624,126 55,455,249 56,409,277 57,387,470 57,829,909 68,213,386 68,233,133 68,245,257 68,296,277 68,307,432 68,429,171 68,395,506
15 54,367,150 53,970,143 55,279,001 55,940,451 57,010,866 57,759,141 58,109,891 68,177,596 68,207,430 68,237,580 68,241,325 68,279,727 68,326,632 68,368,250
16 53,451,225 53,327,809 54,393,699 55,350,104 56,364,900 57,290,723 57,711,596 68,022,941 68,094,829 68,169,848 68,158,831 68,224,797 68,273,585 68,255,991
17 53,645,203 53,634,633 54,177,271 55,078,600 56,138,436 57,251,531 57,541,351 68,138,549 68,147,333 68,221,638 68,287,831 68,325,084 68,415,608 68,411,921
18 54,668,363 54,174,847 55,256,176 56,064,089 56,740,259 57,769,268 57,973,134 68,224,597 68,282,517 68,395,059 68,437,841 68,522,898 68,598,086 68,599,844
19 55,562,101 54,791,623 55,294,695 55,928,589 56,768,136 57,599,180 58,081,287 68,209,538 68,324,444 68,446,514 68,525,618 68,628,926 68,686,765 68,753,261
20 55,938,657 55,081,208 55,877,542 56,253,266 57,048,367 57,967,318 58,106,207 68,136,189 68,313,557 68,470,777 68,543,198 68,639,292 68,722,909 68,829,286
21 56,331,087 55,229,266 55,652,638 56,067,216 56,679,203 57,502,276 57,635,297 68,141,999 68,361,062 68,547,360 68,644,264 68,783,591 68,859,312 68,950,347
22 56,997,804 55,715,430 56,053,895 56,356,954 56,761,453 57,545,255 57,751,704 68,198,529 68,480,921 68,670,594 68,818,205 69,044,454 69,163,424 69,194,391
23 56,802,977 55,429,313 55,735,965 56,197,410 56,471,773 57,057,789 57,383,903 68,143,528 68,525,019 68,831,257 68,946,803 69,230,759 69,321,426 69,445,410
24 57,826,178 56,183,113 56,044,365 55,945,228 55,984,317 56,618,627 56,771,473 68,271,377 68,728,185 69,037,189 69,197,403 69,488,552 69,616,627 69,731,068
25 58,617,239 57,079,668 56,202,206 56,102,928 55,885,915 56,296,069 56,349,525 68,384,381 68,871,654 69,255,113 69,446,599 69,840,923 69,999,312 70,097,498
26 59,309,065 57,207,648 56,075,430 55,816,944 55,552,723 55,814,962 55,846,995 68,608,814 69,166,616 69,653,434 69,885,998 70,350,797 70,565,524 70,654,203
27 59,214,372 56,581,831 55,453,438 55,266,849 54,772,970 55,091,733 55,084,942 68,827,806 69,591,707 70,264,503 70,708,424 71,312,980 71,558,722 71,777,086
28 58,905,505 56,371,887 55,174,878 54,389,172 54,041,511 54,335,228 54,268,336 68,505,025 69,619,124 70,499,723 71,115,782 71,913,592 72,172,923 72,454,327
29 58,023,696 56,122,492 55,053,358 53,998,472 53,301,028 53,274,003 53,087,300 68,732,220 70,492,616 71,644,050 72,524,388 73,553,428 73,967,910 74,302,156
30 56,113,860 54,313,900 52,636,358 51,531,025 50,770,809 50,519,079 50,314,047 74,810,707 77,539,372 79,590,655 80,895,889 82,592,009 83,417,356 83,993,892
31 51,887,230 50,597,329 48,600,945 47,771,643 46,600,073 46,175,796 45,915,086 77,719,313 81,332,604 84,241,907 86,088,401 88,563,161 89,838,828 90,658,855
Avg Context 56,013,884 55,119,707 55,541,817 55,854,922 56,415,273 57,038,583 57,255,126 68,698,135 69,103,111 69,436,860 69,657,280 69,942,103 70,098,195 70,196,268
32 55,823,256 53,884,097 51,681,052 50,569,791 49,394,788 48,863,702 48,537,557 13,930,348 15,076,103 16,016,778 16,465,745 16,788,294 17,279,969 17,455,400
33 58,738,344 56,674,056 54,054,542 52,888,037 51,528,984 50,887,306 50,414,778 26,816,835 27,093,410 27,773,603 27,962,715 28,191,993 28,477,352 28,654,271
34 56,938,764 55,281,907 52,909,883 51,461,738 50,038,303 49,641,208 49,084,752 47,564,034 45,235,994 44,031,407 43,883,655 43,877,411 43,554,253 43,485,497
35 54,366,396 53,029,014 50,888,507 49,755,655 48,770,433 48,334,318 47,981,902 53,101,160 51,120,417 49,742,735 49,185,432 48,518,180 48,081,379 47,894,499
36 52,936,332 51,682,636 49,570,706 48,713,389 47,794,113 47,573,073 47,233,332 58,003,949 56,939,073 55,963,819 55,597,113 55,051,241 54,774,326 54,588,795
37 51,647,223 50,371,630 48,303,608 47,632,078 46,858,461 46,786,053 46,376,421 61,213,159 60,453,961 59,829,451 59,536,354 59,153,093 59,031,495 58,838,218
38 50,931,788 49,209,356 47,576,761 46,731,794 45,968,874 45,790,581 45,477,362 62,655,390 62,137,314 61,620,255 61,458,679 61,231,410 61,110,159 61,003,516
39 51,042,601 49,027,546 47,618,052 46,768,291 45,998,476 45,883,841 45,473,705 63,348,201 62,998,381 62,683,843 62,649,981 62,440,555 62,378,970 62,319,582
40 49,590,360 47,756,219 46,431,515 45,965,206 45,449,164 45,252,564 45,184,300 63,866,857 63,600,670 63,423,587 63,430,422 63,301,531 63,282,617 63,207,398
41 49,954,374 47,928,574 46,848,807 46,364,193 45,679,210 45,474,688 45,440,294 64,301,293 64,110,817 63,913,386 63,893,724 63,873,160 63,880,434 63,857,697
42 50,897,285 48,822,840 47,723,273 47,389,348 46,461,956 46,336,541 46,106,581 64,486,391 64,515,591 64,453,381 64,379,111 64,406,903 64,442,057 64,385,303
43 52,467,447 49,948,494 48,687,532 48,087,823 47,426,244 47,134,027 46,970,169 64,557,953 64,566,576 64,532,662 64,488,626 64,565,552 64,573,767 64,585,706
44 54,876,822 52,145,505 50,513,749 49,816,893 48,819,838 48,428,789 48,281,333 64,662,055 64,739,419 64,747,685 64,807,289 64,861,456 64,982,075 64,897,827
45 56,056,772 53,288,677 51,530,428 50,709,794 49,538,754 49,171,650 48,847,311 64,772,578 64,907,757 64,948,826 64,997,734 65,068,658 65,124,986 65,132,858
46 61,167,369 59,469,476 57,204,341 55,850,490 54,489,360 53,898,637 53,288,932 64,727,227 64,840,307 64,926,333 64,908,159 65,053,638 65,091,642 65,154,625
47 71,009,467 67,839,176 66,565,545 65,503,549 63,999,825 63,292,284 62,803,150 64,590,446 64,867,523 64,840,482 64,903,198 65,014,873 65,036,948 65,120,427
48 66,672,308 64,419,151 63,096,912 62,676,497 61,761,157 60,991,890 60,750,602 66,612,523 66,803,144 66,905,353 66,952,648 67,075,374 67,116,770 67,266,114
Avg Continuation 55,595,112 53,575,197 51,835,601 50,993,210 49,998,702 49,631,832 49,308,969 57,012,377 56,706,262 56,491,387 56,441,211 56,380,784 56,365,835 56,343,984
Avg Sentence 55,865,569 54,572,693 54,229,199 54,133,066 54,142,738 54,415,359 54,440,862 64,559,429 64,712,561 64,852,005 64,976,589 65,139,135 65,234,651 65,290,251
Table 5: 2 Gram Statistics for memorized and unmemorized content
target half half half half half half half quarter quarter quarter quarter quarter quarter quarter
size 70m 160m 410m 1b 2.8b 6.9b 12b 70m 160m 410m 1b 2.8b 6.9b 12b
1 53,888,945 55,774,256 56,835,849 57,399,791 58,749,349 58,746,839 59,572,656 65,365,004 66,059,083 66,813,871 67,057,786 67,482,982 67,249,886 67,522,493
2 53,633,778 55,585,291 56,294,740 57,125,864 58,375,426 58,787,083 59,024,756 65,505,804 66,197,266 66,585,597 66,774,723 67,277,478 67,415,528 67,313,805
3 53,966,515 55,464,751 56,571,797 57,248,585 58,522,726 59,007,609 59,282,618 65,526,925 66,301,214 66,609,558 67,101,450 67,242,630 67,240,433 67,479,250
4 53,788,979 55,012,249 55,852,524 56,732,428 58,632,577 58,585,481 59,106,526 65,575,121 65,977,326 66,498,055 66,914,635 67,144,082 67,128,936 67,304,810
5 52,887,357 54,130,559 55,380,593 56,487,949 57,789,441 58,410,016 58,836,366 65,413,320 66,086,522 66,640,160 66,821,619 67,103,545 67,257,537 67,368,841
6 53,346,585 54,626,718 56,348,320 56,603,397 58,159,466 58,648,425 58,945,615 65,113,237 66,153,641 66,499,304 66,682,733 67,063,613 67,252,679 67,384,155
7 52,836,947 54,888,957 55,457,422 56,613,020 57,855,215 58,349,009 58,580,014 65,090,106 65,649,597 66,459,689 66,678,897 66,947,213 67,002,958 67,127,891
8 52,478,474 54,201,487 55,628,345 56,114,532 57,493,118 58,018,440 58,920,964 64,824,999 65,498,431 66,241,656 66,399,994 66,746,465 66,760,013 66,819,417
9 53,600,100 55,410,940 56,362,313 57,618,326 58,256,347 59,088,717 59,814,719 64,941,905 65,788,899 66,460,344 66,506,701 66,921,900 67,080,681 67,174,870
10 53,868,058 55,087,199 56,362,786 57,058,982 58,314,177 58,745,840 59,526,404 64,563,380 65,502,285 66,335,635 66,276,258 66,961,511 66,922,326 66,965,502
11 52,884,280 54,383,701 56,075,362 56,818,188 58,050,191 58,352,813 58,596,784 64,715,807 65,616,229 66,029,659 66,452,694 66,712,759 66,674,522 66,995,794
12 52,928,646 54,849,387 55,982,049 56,771,448 58,094,228 58,565,819 58,811,948 64,820,938 65,424,995 65,790,449 66,253,435 66,525,526 66,698,298 66,719,967
13 52,690,921 54,559,756 55,697,194 56,532,692 57,846,992 58,355,202 58,946,804 64,842,590 65,142,936 65,893,506 66,065,102 66,465,922 66,738,926 66,653,765
14 52,781,679 54,368,904 56,019,791 56,847,208 57,869,958 58,242,828 58,685,929 64,404,765 65,245,584 65,797,393 66,052,000 66,479,573 66,541,779 66,586,451
15 53,765,298 55,537,152 56,289,403 57,279,213 58,113,960 58,741,123 59,494,649 64,231,835 64,970,624 65,603,603 65,691,554 66,258,007 66,393,008 66,421,189
16 53,740,078 55,037,415 55,929,818 56,941,982 57,702,245 58,282,107 58,697,163 64,285,467 64,973,050 65,644,470 65,985,554 66,366,943 66,449,639 66,680,453
17 53,610,921 55,130,291 56,781,859 57,648,192 58,161,184 58,670,246 59,044,372 64,584,535 65,324,214 66,141,495 66,173,896 66,423,703 66,634,663 66,523,072
18 52,403,931 54,507,436 56,143,121 56,609,104 57,872,099 58,070,077 58,324,067 64,823,801 65,426,701 65,725,421 65,794,806 66,079,791 66,355,245 66,570,526
19 51,734,926 53,970,664 55,420,133 56,429,139 57,389,684 58,026,044 57,807,255 64,435,084 65,175,915 65,608,367 65,736,235 66,057,197 66,188,760 66,351,043
20 52,066,233 53,844,254 55,306,358 56,168,419 56,834,780 57,322,399 57,930,955 64,026,114 64,482,213 64,845,907 65,324,118 65,663,196 65,946,061 65,795,199
21 54,499,372 55,151,054 55,817,487 56,104,835 57,282,099 57,680,431 58,491,658 63,854,665 64,377,391 64,858,096 64,933,132 65,479,905 65,461,356 65,605,633
22 53,002,234 53,689,249 54,689,980 55,709,870 56,393,115 56,890,668 57,461,711 63,953,011 64,210,926 64,748,681 64,882,441 65,007,446 65,115,475 65,219,898
23 54,475,618 54,209,390 55,114,979 55,800,549 56,745,566 56,890,699 57,313,138 63,693,211 64,323,138 64,361,084 64,798,794 64,818,666 65,024,302 64,992,457
24 52,696,844 53,885,907 54,088,030 55,057,040 55,463,131 56,208,303 56,444,958 63,400,561 63,820,771 63,700,022 64,438,090 64,258,752 64,141,101 64,186,891
25 50,318,283 51,269,205 52,481,347 53,465,984 53,891,662 54,781,920 54,713,197 63,232,652 63,129,362 63,259,586 63,725,082 63,910,320 63,867,932 63,988,413
26 50,229,788 51,354,758 52,362,605 53,242,078 53,893,195 54,326,699 54,530,116 61,826,182 61,893,288 62,274,584 62,752,607 62,746,811 62,887,121 63,116,564
27 52,206,436 52,688,837 53,341,050 53,769,840 54,270,899 54,576,546 55,060,048 60,352,198 60,726,965 60,923,515 61,131,020 60,963,222 61,265,266 61,663,726
28 49,682,706 50,249,019 51,876,101 52,495,005 52,542,959 53,202,973 53,174,103 60,534,641 60,288,939 60,079,857 59,975,267 60,096,463 59,981,804 60,079,474
29 49,857,350 50,055,633 51,690,368 51,477,619 51,245,707 51,586,655 52,023,933 61,446,881 60,521,500 59,891,967 59,610,978 59,336,539 59,242,416 59,109,188
30 48,055,193 47,429,451 48,033,530 48,151,423 48,174,882 48,183,182 48,215,800 59,390,999 57,149,284 56,340,637 55,831,719 55,327,710 54,997,345 54,953,972
31 39,251,259 39,137,018 39,388,512 39,273,372 39,230,840 39,250,211 39,423,466 47,469,654 45,842,988 45,103,473 44,628,970 44,396,335 44,093,542 44,099,541
Avg Context 52,167,024 53,402,932 54,503,993 55,212,777 56,103,781 56,535,303 56,929,119 63,427,271 63,783,267 64,121,472 64,304,913 64,524,716 64,580,953 64,670,137
32 50,487,094 49,193,302 48,748,380 48,353,375 48,143,528 47,788,231 47,682,812 73,626,535 69,054,469 66,707,831 64,778,405 63,608,111 62,438,327 62,279,482
33 61,392,551 59,894,701 58,602,818 58,523,094 57,350,370 57,349,346 56,865,365 127,465,382 117,081,737 110,181,057 106,211,754 102,176,536 100,298,529 98,864,531
34 61,328,405 61,880,320 61,897,331 61,603,725 61,037,948 60,910,761 61,148,637 183,840,986 178,535,599 174,318,924 169,430,877 164,785,135 162,605,626 161,446,058
35 62,108,896 64,859,473 66,373,167 66,487,737 66,963,760 67,607,923 67,185,378 207,065,353 212,630,432 212,924,641 212,313,484 210,389,783 209,373,568 209,264,567
36 64,491,225 66,128,068 68,573,633 68,479,538 69,624,085 69,468,930 70,076,197 107,042,066 102,720,927 98,446,681 95,544,431 91,884,559 89,921,371 89,176,573
37 64,176,596 67,738,610 71,189,574 72,648,126 74,649,180 75,348,320 76,011,136 78,418,713 75,370,613 73,491,350 72,095,517 70,419,121 69,798,088 69,480,522
38 83,023,501 86,268,700 92,242,626 95,541,179 100,137,087 103,141,517 104,996,137 75,331,828 73,633,103 72,372,719 71,451,162 70,344,087 69,870,019 69,555,251
39 114,997,010 123,345,929 132,119,453 137,345,936 144,729,841 148,911,774 152,506,274 72,416,590 71,374,676 70,711,375 69,920,343 69,170,308 68,691,759 68,547,626
40 72,923,797 75,060,147 76,865,835 77,568,398 78,566,453 79,311,885 80,082,621 73,411,371 72,849,906 71,665,061 71,318,529 70,698,319 70,306,675 70,168,524
41 62,139,748 64,712,748 65,532,787 65,890,626 66,512,689 66,430,631 67,029,855 73,147,695 72,602,565 71,681,923 71,271,172 70,741,320 70,453,928 70,443,127
42 61,518,278 63,119,965 63,331,920 63,642,209 64,136,809 63,861,430 63,990,245 72,618,045 71,781,987 71,440,631 71,063,281 70,644,161 70,394,990 70,299,681
43 58,893,899 59,747,917 60,630,537 60,111,792 60,897,906 60,413,416 60,685,337 72,647,660 71,716,108 71,179,077 70,986,376 70,431,462 70,058,904 70,263,439
44 57,229,696 57,312,370 57,433,177 58,104,220 57,890,592 58,313,878 58,439,482 72,211,146 71,314,701 70,977,086 70,897,382 70,417,594 70,322,615 70,408,343
45 55,873,454 56,139,949 56,811,592 56,589,596 57,622,149 57,870,333 58,004,104 72,853,238 71,729,416 71,312,109 71,062,683 70,958,396 70,941,043 70,812,855
46 54,966,654 55,910,428 56,508,715 56,557,283 57,408,451 57,383,704 58,025,074 73,234,042 72,148,832 71,787,141 71,388,367 71,441,543 70,955,811 71,090,486
47 55,258,100 55,540,073 56,121,577 56,868,270 57,182,726 57,490,540 58,014,824 71,452,519 70,861,810 70,783,165 70,262,176 70,511,868 70,420,548 70,296,965
48 53,806,444 54,635,210 54,818,474 55,121,630 55,692,510 56,391,570 56,482,717 65,034,377 65,452,587 65,749,667 66,046,086 66,391,820 66,416,952 66,424,818
Avg Continuation 64,389,138 65,969,877 67,517,741 68,202,161 69,326,240 69,882,011 70,425,070 92,459,856 90,638,792 89,160,614 88,002,472 86,765,537 86,074,632 85,813,109
Avg Sentence 56,495,689 57,853,725 59,113,028 59,813,184 60,786,735 61,262,262 61,708,935 73,709,645 73,294,599 72,989,502 72,697,798 72,401,674 72,193,298 72,158,273
Table 6: 2 Gram Statistics for half-memorized and quarter-memorized content.